### v28

# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

Abstract
We analyze price discovery in a model where an agent has arbitrary private information about state probabilities and trades state-contingent claims. Our model integrates the seminal frameworks of Arrow and Debreu (1954) and Kyle (1985). In an equivalent options formulation, the informed agent has arbitrary information about an underlying asset’s payoff distribution and trades option portfolios. We characterize the informed demand and information dynamics across markets. Our results offer the first equilibrium-based explanations for longstanding empirical practices and phenomena in option markets, including common trading patterns and the volatility smile across option strikes.
JEL classification: C61; G13; G14.
Keywords: Price Discovery, Contingent Claims, Options, Straddle, Volatility Smile.
1 Introduction
A fundamental economic function of prices is to convey information. As informed agents trade to exploit their information about an asset’s expected payoff, that information is incorporated into the asset price. More broadly, agents often have information about an asset’s payoff beyond its mean, such as volatility, and to leverage such information they must trade derivatives. Indeed, a principal utility of derivatives is to offer investors targeted exposure to various aspects of the underlying asset’s payoff distribution. Such derivatives trading has long been standard practice, as covered in popular textbooks such as Hull (2015). This implies that a wide range of information beyond just the expected payoff is incorporated into derivatives prices.
Yet, despite their practical importance, the existing literature lacks a basic model that articulates how derivative markets systematically incorporate diverse information signals into prices. This gap stems largely from the challenges of modeling the complex information dynamics across derivatives. We address this gap by developing a model that fully captures the nonlinearity of payoff distributions and the interdependence of trades and prices across derivatives. Integrating the foundational frameworks of Arrow and Debreu (1954) and Kyle (1985), we obtain general characterizations of the informed demand, price impact, and information efficiency of prices across derivative markets.
In our model, an informed trader has private information about the probabilities of future states and trades state-contingent claims. In other words, the informed trader has information about the payoffs of Arrow-Debreu (AD) securities and trades in the AD markets. There is no restriction on the probability distribution across states or the type of information available to the informed trader.
Replicating derivatives, such as options, using AD securities provides equivalent formulations of our model that account for the structural linkages across derivatives. In an equivalent options formulation, the informed trader has information about an underlying asset’s payoff distribution and trades options, with no restriction on the payoff distribution or his information.
The informed demand formula we obtain is simple and practical. It constructs the informed trader’s contingent-claim portfolio by assigning positive portfolio weight to the true payoff distribution and negative weights to the others. Translated to options, the formula prescribes option strategies for trading on any given aspect of the underlying payoff distribution. In particular, it offers the first equilibrium rationalization for widely used option trading strategies.
When trading is restricted to only the underlying asset, our model reduces to the single-asset price discovery framework of Kyle (1985), where the equilibrium supply function is linear. The slope of the inverse supply function, known as “Kyle’s lambda,” determines the price impact and the extent to which demand reveals information about the asset’s payoff. This slope is proportional to the variation of the asset’s payoff. In our more general AD setting, demand for one security impacts prices across all markets based on the information it conveys about other securities. We show that the equilibrium price impact between two securities is proportional to their payoff covariance, generalizing Kyle’s lambda to an infinite-dimensional matrix of payoff covariances.
Across option markets, trades in one option often impact the prices of others as options trades jointly convey information about the underlying payoff. For example, a common volatility trade involves trading a put-call pair to profit from price movements regardless of direction. Therefore, trades in such pairs jointly convey volatility information. Our characterization of such cross-market information spillovers generates new empirical predictions linking cross-option dynamics to various aspects of the underlying payoff. Moreover, we offer an equilibrium explanation for the “volatility smile,” an empirical regularity of option-implied volatilities that has been widely studied but only in reduced-form.
The joint supply schedules for all AD securities form a pricing kernel, under asymmetric information. In contrast to the one-dimensional linear supply function for a single asset, this pricing kernel is a high-dimensional nonlinear object that prices all derivatives simultaneously. While the pricing of derivatives has been extensively studied from the hedging perspective since Black and Scholes (1973), we go beyond hedging under complete information, providing a microfoundation for derivatives trading and price patterns that hedging motives alone cannot explain.
While empirical evidence (e.g., Roll et al. (2010)) suggests derivatives improve market efficiency, theoretical explanations are limited. Our model offers a clear explanation: the linear equilibrium of Kyle (1985) is not robust with respect to the introduction of derivatives, as the absence of these markets constrains the informed trader to suboptimal linear demands. Derivatives allow the informed trader to optimally reduce his demand for low-probability states (see Example 7.1). By incorporating information that cannot be reflected in linear trades of the underlying asset, derivatives improve market efficiency.
In a complete market with derivatives, our model demonstrates that the information efficiency of all prices is summarized by a single measure. Notably, this measure remains invariant to variations in the possible payoff distributions or noise trading levels across markets. Thus, the informativeness of derivatives prices is robust to specific information details.
Related Literature
Kyle (1985) provides the seminal framework for single-asset price discovery. Extensions by Caballe and Krishnan (1994), Foster and Viswanathan (1996), and Back et al. (2000) consider settings with multiple insiders holding differential information across assets. Rochet and Vila (1994) relax the assumption of normal payoff distributions. We expand the scope of this literature to encompass joint price discovery across contingent claim markets.
Empirical research has consistently shown that trades and prices across options markets reflect information beyond the expected payoff. Pan and Poteshman (2006) and Ni et al. (2008) find that option trading volumes predict future stock returns and volatility, while Goyal and Saretto (2009) show that volatility mispricing can be exploited through option straddles. Other studies highlight how option order flows can signal firm value and market expectations of future volatility—see Roll et al. (2009), Augustin et al. (2019), and Cao et al. (2005). Despite substantial empirical evidence of joint price formation across option markets, existing theoretical models remain limited to no-option versus single-option comparisons (e.g., Back (1993); Biais and Hillion (1994); Easley et al. (1998); Collin-Dufresne et al. (2021)). This leaves an empirical-theoretical gap in understanding how trades and prices across options jointly reflect broader information. Our model bridges this gap by directly capturing the richness of information embedded across option trades and prices. It provides a unified theoretical framework that consolidates empirical findings and suggests new avenues of empirical research.
Different from our strategic setting, there have also been theoretical studies on non-strategic informed trading across multiple securities, such as Admati (1985), Malamud (2015), and Chabakauri et al. (2022). These non-strategic models do not explain the common derivatives trading strategies or observed cross-strike empirical regularities. Moreover, Admati (1985) and Chabakauri et al. (2022) impose restrictive parametric assumptions on the payoff distribution, and Malamud (2015) depends on properties intrinsic to the continuum. We make no parametric assumptions on the payoff, and our results do not depend on the discrete or continuum modeling choice. Additionally, we derive testable predictions linking the underlying payoff characteristics, such as higher moments, to observed price impacts across options. This new insight is unique to the strategic setting.
Many option pricing and financial econometric models have sought to capture the volatility smile by correcting misspecifications in the Black and Scholes (1973) framework. Prominent classes of such models include the jump-diffusion, [^1] See Merton (1976) and Bates (1991). stochastic volatility, [^2] See Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), and Aït-Sahalia et al. (2021). and local volatility models. [^3] See Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012). Our framework moves beyond reduced-form model misspecifications to directly explain the strike structure of implied volatility via cross-strike information dynamics.
The rest of the paper proceeds as follows. Section 2 demonstrates the basic intuition of our results through a simplified setting. Section 3 presents the general model. Section 4 identifies the sufficient statistic for the uninformed trader’s cross-market inferences. Section 5 analyzes the informed trader’s cross-market price impact. Section 6 establishes equilibrium through a canonical reduction of the general trading game. While this canonical reduction is a key modeling innovation that makes our general model tractable, the broad intuition of the resulting equilibrium extends directly from Section 2 and can be understood from Example 6.7 without delving into the precise details. Section 7 discusses the characterization of price discovery across contingent claims—the informed demand, price impact, and information efficiency of prices. Section 8 explores the immediate empirical implications of our results for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
(a) $x$-axis: states, $y$-axis: probability
(b) $x$-axis: insider portfolio size $\alpha$, $y$-axis: expected profit
Figure 1: Simplified Setting
(a) Conditional State Probabilities. Solid dots represent signal $s_{1}$ (high volatility), and hollow dots represent signal $s_{2}$ (low volatility).
Insider Expected Profit Curve. The insider’s optimal portfolio size $\alpha^{*}$ characterizes equilibrium.
2 Basic Intuition
We begin with a simplified setting that encapsulates the basic economic intuition we bring out. In the general model, we will develop the robustness of this intuition—without restriction on the number of securities, payoff distributions, or type of private information.
There are two risk-neutral agents: the insider and the market maker. At $t=0$, the agents know that there are three possible $t=1$ states, denoted $x_{i}$ for $i=1,2,3$, and trade the corresponding AD securities.
At $t=0$, the insider privately observes one of two possible signals, $s_{1}$ or $s_{2}$. Conditional on signal $s_{1}$, the probability distribution over $t=1$ states is $(\frac{1}{2},0,\frac{1}{2})$. Conditional on $s_{2}$, the distribution is $(0,1,0)$. See Figure 1a. The market maker has the uniform prior $\pi_{0}$ on signals.
After observing his private signal, the insider submits his demand (market orders) for AD securities to maximize his expected utility at $t=1$. Noise traders trade for exogenous reasons such as liquidity needs, and their trades across the AD markets are $\varepsilon_{i}\!\!\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\!\!\mathcal{N}(0,1)$ for $i=1,2,3$. The market maker acts as a competitive liquidity provider.
[…]
$(P_{1},P_{2},P_{3})=(\frac{1}{2}\pi_{1}(s_{1}|\omega),\,\pi_{1}(s_{2}|\omega),\,\frac{1}{2}\pi_{1}(s_{1}|\omega)).$
An intuitive trading strategy for the insider, conditional on observing $s_{1}$, is to buy securities $x_{1}$ and $x_{3}$ (where payoffs are positive) and sell security $x_{2}$ (where payoff is zero) for some equal $\alpha$ shares each. [^5] There are no leverage or short-selling constraints. Then, conditional on the insider observing $s_{1}$, the order flow across markets received by the market maker is
$\omega=(\omega_{1},\omega_{2},\omega_{3}),\;\mbox{ where }\;\omega_{1}=\alpha+\varepsilon_{1},\;\omega_{2}=-\alpha+\varepsilon_{2},\;\omega_{3}=\alpha+\varepsilon_{3}.$
Similarly, conditional on $s_{2}$, the insider buys security $x_{2}$ and sells securities $x_{1}$ and $x_{3}$ for $\alpha$ shares each.
Market Maker’s Sufficient Statistic
Given the insider’s trading strategy, the market maker’s posterior likelihood ratio, conditional on order flow $\omega$, is
$\frac{\pi_{1}(s_{1}|\omega)}{\pi_{1}(s_{2}|\omega)}=\frac{{\rm e}^{\alpha\omega_{1}-\frac{1}{2}\alpha^{2}-\alpha\omega_{2}-\frac{1}{2}\alpha^{2}+\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}}{{\rm e}^{-\alpha\omega_{1}-\frac{1}{2}\alpha^{2}+\alpha\omega_{2}-\frac{1}{2}\alpha^{2}-\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}}.$ (1)
Therefore, his zero-profit AD prices are
$(P_{1},P_{2},P_{3})=(\frac{\frac{1}{2}{\rm e}^{\alpha\Delta}}{{\rm e}^{\alpha\Delta}+1},\,\frac{1}{{\rm e}^{\alpha\Delta}+1},\,\frac{\frac{1}{2}{\rm e}^{\alpha\Delta}}{{\rm e}^{\alpha\Delta}+1}),\quad\mbox{ where }\;\Delta(\omega)=2(\omega_{1}+\omega_{3}-\omega_{2}).$ (2)
Thus, the market maker’s sufficient statistic is the excess demand $\Delta(\omega)$ for $x_{1}$ and $x_{3}$ securities relative to $x_{2}$. If $\Delta(\omega)$ is high, he infers the insider likely observed $s_{1}$, leading him to raise $P_{1}$ and $P_{3}$ and lower $P_{2}$, and vice versa. In other words, there is cross price impact, and it is determined by the covariances of security payoffs. Since the payoff covariance of securities $x_{1}$ and $x_{3}$ is $\frac{1}{8}$, their cross price impact is positive. The pairwise payoff covariance for the other two security pairs is $-\frac{1}{8}$, resulting in negative cross price impact. The payoff covariances determine the extent to which order flow for one security reveals information about others, and thus determine the price impact across markets. This intuition extends directly to the general model developed in subsequent sections.
Insider Portfolio Choice
The insider, conditional on, say, $s_{1}$, scales his buy-sell (long-short) portfolio by $\alpha$ to maximize expected profit
$\max_{\alpha\geq 0}\,\mathbb{E}[\alpha\,(\frac{1}{2}+\frac{1}{2}-P_{1}+P_{2}-P_{3})].$
Figure 1b illustrates the insider’s expected profit as a function of $\alpha$. If $\alpha\!=\!0$, the insider’s zero demand results in zero profit, and the AD prices are $(\frac{1}{4},\frac{1}{2},\frac{1}{4})$; since order flow contains no information, the market maker sets the prices based on his prior. If $\alpha\!\rightarrow\!\infty$, the market maker’s posterior probability of $s_{1}$ approaches one, which causes the AD prices to converge to the true distribution $(\frac{1}{2},0,\frac{1}{2})$, driving the insider’s profit to zero due to full information revelation.
The insider’s optimal portfolio size $\alpha^{*}$ trades off the marginal payoff from exploiting the true distribution $(\frac{1}{2},0,\frac{1}{2})$ against the marginal costs. These costs include the cost incurred at current prices, assuming no price impact, and the cost stemming from the cross-market price impact due to information revelation. His optimal portfolio size $\alpha^{*}$ characterizes equilibrium—this simple result will extend to the general model. Allowing correlated noise trades, varying noise intensity, and non-uniform priors does not materially affect the results. [^6] Noise Trades Correlation Allowing noise trades to be correlated, while keeping their intensity constant, does not affect price impact. For example, suppose the correlation of noise trades for a pair of assets has an opposite sign to their payoff correlation. In that case, the insider would scale up his trading strategy to capture additional losses incurred by the noise traders trading against the payoffs. The market maker would shift his supply schedules accordingly, but the price impact at the margin would remain the same; this is similar to biased noise trades in the single-asset setting.  [^7] Noise Trading Intensity Suppose the noise trading intensity $(\sigma_{i})$ varies across markets, with $\epsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$, $i=1,2,3$. Then $\Delta(\omega)$ becomes $\Delta(\omega)=2(\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}).$ Thus, as expected, higher noise trading intensity reduces price impact within each market and across markets.  [^8] General Prior When the market maker has a general prior $(\pi_{0}(s_{1}),\pi_{0}(s_{2}))$, the posterior likelihood ratio (1) becomes $\frac{\pi_{1}(s_{1}|\omega)\pi_{0}(s_{1})}{\pi_{1}(s_{2}|\omega)\pi_{0}(s_{2})}=\frac{e^{\alpha\omega_{1}-\frac{1}{2}\alpha^{2}-\alpha\omega_{2}-\frac{1}{2}\alpha^{2}+\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{1})}{e^{-\alpha\omega_{1}-\frac{1}{2}\alpha^{2}+\alpha\omega_{2}-\frac{1}{2}\alpha^{2}-\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{2})}.$ The comparative statics for the prior follows intuitively. For example, a higher $\pi_{0}(s_{1})$ leads to greater cross price impact between securities $x_{1}$ and $x_{3}$—because it leads the market maker to place more emphasis on the co-movement of the orders for $x_{1}$ and $x_{3}$ in his inference.
Information Role of Derivatives
If fewer than all three AD securities are available for trade, both the insider’s profit and the information efficiency of prices decrease. Here, information efficiency is naturally measured by the posterior probability weight that the market maker assigns to the true signal. Derivative markets allow the informed agents’ trades to better reflect their information, thus incorporating their private information into prices. This is the basic economic logic motivating this paper.
Example: “Straddle”
Suppose the states $x_{1}<x_{2}<x_{3}$ are the possible $t=1$ prices of an underlying asset whose $t=0$ price is $x_{2}$. In this case, $s_{1}$ is the “high volatility” signal, and $s_{2}$ the “low volatility” signal. The signal $s_{1}$ reveals that the asset price will move, but without a specified direction, whereas $s_{2}$ reveals that there will be no price movement. Security $x_{1}$ acts as a toy put option that pays off when the asset price declines, while $x_{3}$ acts as a call option that pays off when the asset price rises. If the insider observes the high volatility signal $s_{1}$, he buys a toy put-call pair—a straddle—to capitalize on price movement regardless of direction. The actual straddle is a standard option strategy for volatility trading (see Hull (2015)) and is widely discussed in the empirical literature. [^9] See Coval and Shumway (2001), Pan and Poteshman (2006), Ni et al. (2008), Driessen et al. (2009), and Goyal and Saretto (2009). This strategy, along with other common option strategies, will be explained as equilibrium strategies in the general model.
The cross price impact between the put-call pair in the toy straddle is greater under the high volatility signal, because the insider’s demand for the pair amplifies their cross-market information spillover. Therefore, information about the future volatility of the underlying payoff can be backed out from the cross price impact between the put-call pair in the straddle. In our general model, we systematically extend this insight to actual options and derive empirical predictions about general aspects of the underlying asset’s payoff.
3 General Model
As in Section 2, there are two risk-neutral agents, the insider and the market maker. At $t=0$, the insider observes a signal that informs him of the probability distribution over the set of possible $t=1$ states, $X=\{x_{1},x_{2},\cdots\}$ (which is countably infinite, with finite $X$ included as a special case). The market maker has a Bayesian prior over the possible signals. At $t=0$, there is a complete market of AD securities for $t=1$ states. [^10] The model can be adapted to the incomplete market case by partitioning the state space. In this paper, we focus on the complete market case. There is a risk-free asset in perfectly elastic supply at risk-free rate zero.
After observing his private signal at $t=0$, the insider submits his demand for AD securities to maximize his expected utility at $t=1$. The market maker then receives the combined order flow of the insider and noise traders across AD markets and executes the orders at his zero-profit prices.
The set of possible signals is $S=\{s_{1},\cdots,s_{K}\}$, with the market maker holding prior $\pi_{0}(s_{k})$, $1\leq k\leq K$. Conditional on a signal $s_{k}$, the probabilities of $t=1$ states are given by $\eta(\,\cdot\,|s_{k})\colon X\rightarrow[0,\infty)$. After observing $s_{k}$, the insider chooses a portfolio $W(\,\cdot\,|s_{k})\colon X\rightarrow\mathbb{R}$ where $W(x_{i}|s_{k})$ is his order for security $x_{i}$ conditional on $s_{k}$.
Noise trader orders $\varepsilon_{i}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,\sigma_{i}^{2})$, $i\geq 1$, are normally distributed with mean zero in each market and uncorrelated across markets. When the insider chooses portfolio $W\colon X\rightarrow\mathbb{R}$, the order flow received by the market maker is a stochastic sequence $\omega=(\omega_{i})$ where
$\omega_{i}=W(x_{i})+\varepsilon_{i}\quad\mbox{for each market $i$.}$
The market maker holds a belief $\widetilde{W}(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$ about the insider’s trading strategy. Based on this belief and the received order flow $\omega$, the market maker updates his prior $\pi_{0}$ to the posterior $\pi_{1}(s_{k}|\omega\kern 0.44998pt;\kern 0.44998pt\widetilde{W})$, $1\leq k\leq K$, over signals. His zero-profit price for each security $x_{i}$ is his posterior mean of its payoff $\eta(x_{i}|\,\cdot\,)$,
$\underbrace{P(x_{i}|\omega\kern 0.44998pt;\widetilde{W})}_{\text{security $x_{i}$ price}}=\sum_{k}\!\eta(x_{i}|s_{k})\,\pi_{1}(s_{k}|\omega\kern 0.44998pt;\kern 0.44998pt\widetilde{W})\quad\mbox{for each market $i$.}$ (3)
Conditional on observing $s_{k}$ and given market maker belief $\widetilde{W}$, the insider’s AD portfolio choice problem is
[…]
Here, the expectation $\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[\,\cdot\,]$ is taken over order flow $\omega$ under its distribution $\mathbb{P}_{\!\scriptscriptstyle W}$ induced by $W$. The functional $J(\,\cdot\,|s_{k};\widetilde{W})$, as defined in (4), is the insider’s expected utility functional conditional on $s_{k}$ and given market maker belief $\widetilde{W}$.
In equilibrium, the insider’s optimal trading strategy given the market maker’s pricing rule $P(\,\cdot\,|\,\cdot\,\kern 0.44998pt;\kern 0.44998ptW^{*})$ based on the latter’s belief $W^{*}$ coincides with $W^{*}$. That is, conditional on each $s_{k}$, the insider’s optimal portfolio is $W^{*}(\,\cdot\,|s_{k})$, thus confirming the market maker’s belief.
Definition 1.
A (Perfect Bayesian) equilibrium in our model is a trading strategy $W^{*}(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$ such that, conditional on each signal $s_{k}$,
$W^{*}(\,\cdot\,|s_{k})\in\operatorname*{argmax}\limits_{W\colon X\rightarrow\mathbb{R}}J(W|s_{k};W^{*}).$
Remark 3.1.
There are no restrictions on $\eta(\,\cdot\,|\,\cdot\,)$, i.e., no restrictions on the probability distributions across states or the insider’s information.
Single-Asset Special Case (Kyle (1985))
Reducing $X$ to a singleton yields the single-asset special case of our model. In this case, the insider’s information $\eta(\,\cdot\,|s):X\rightarrow[0,\infty)$ can be collapsed to just the signal $s$ that parameterizes the asset’s expected payoff at $t=1$. This is the (static) Kyle (1985) setting.
The insider’s strategy reduces to $W\colon S\rightarrow\mathbb{R}$, where he submits a market order $W(s)$ for the asset after observing $s\in S$. When the insider submits order $W$, the market maker receives the combined order $\omega=W+\varepsilon$, where $\varepsilon$ is the noise trader order.
The market maker has a belief $\widetilde{W}\colon S\rightarrow\mathbb{R}$ about the insider’s strategy and, based on this belief, forms his posterior regarding the asset value conditional on $\omega$. His zero-profit price $P(\omega\kern 0.44998pt;\widetilde{W})$ for the asset is his posterior mean. The insider’s problem, conditional on observing $s$ and given market maker belief $\widetilde{W}$, is
$\max_{W\in\mathbb{R}}\,\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[(s-P(\omega\kern 0.44998pt;\widetilde{W}))\cdot W]$
where the expectation $\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[\,\cdot\,]$ is taken over order flow $\omega$ with respect to its distribution $\mathbb{P}_{\!\scriptscriptstyle W}$ induced by his order $W$.
An equilibrium is a trading strategy $W^{*}(\,\cdot\,)$ such that, conditional on each possible asset value $s$,
$W^{*}(s)\in\operatorname*{argmax}_{W\in\mathbb{R}}\,\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[(s-P(\omega\kern 0.44998pt;W^{*}))\cdot W].$
This is Definition 1 of Kyle (1985) and a special case of our Definition 1 when there are no derivative markets. [^11] In Definition 1 of Kyle (1985), his Equation (2.2)—$\tilde{p}(X,P)=E\{\tilde{v}|\tilde{x}+\tilde{u}\}$ where $\tilde{x}=X(\tilde{v})$—means that, in his notation, the insider’s optimal strategy $X(\,\cdot\,)$ confirms the belief that underlies the market maker’s pricing rule $P(\,\cdot\,)$.
Assume the asset value follows a normal prior distribution $\pi_{0}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(v_{0},\sigma_{v}^{2})$ with prior mean $v_{0}$ and variance $\sigma_{v}^{2}$, and that noise trades follow $\varepsilon\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{\varepsilon}^{2})$. Under these assumptions, there exists a linear equilibrium
$W^{*}(s)=\beta(s-v_{0})\,\mbox{ and }\,P(\omega\kern 0.44998pt;W^{*})=v_{0}+\lambda\omega$ (5)
where
$\beta=\frac{\sigma_{\varepsilon}}{\sigma_{v}}\,\mbox{ and ({\em Kyle's lambda}) }\lambda=\frac{\sigma_{v}}{2\sigma_{\varepsilon}}.$ (6)
This is Theorem 1 of Kyle (1985).
Remark 3.2.
Intuitively, the normal prior can be approximated on a discrete grid of signals. All our results hold regardless of whether the state and signal spaces are modeled as discrete or continuous. The economic intuition we bring out is robust to either modeling choice, but the formal proofs for continuous spaces are considerably more technical. We provide the complete mathematical analysis for the continuous case in a separate paper.
In this single-asset special case, the price impact is the slope $\lambda$ of the market maker’s inverse supply function $P(Q)=v_{0}+\lambda Q$, as specified by Equations (5) and (6). Price impact is proportional to the noise-adjusted signal variation $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$. Higher signal variation leads to greater informed demand variation across signals, making order flow more informative for the market maker and resulting in a higher price impact and reduced informed demand. Thus, the price impact and informed demand are determined by the information intensity of the insider’s signal. As suggested in Section 2, this intuition generalizes across AD markets.
Equivalent Options Formulation
Suppose the states $x_{1}<x_{2}<\cdots$ are the possible $t=1$ prices of an underlying asset. In this case, any portfolio $W(\,\cdot\,)$ of AD securities can be replicated using a combination of the underlying asset and options: [^12] See Back (2010) Exercise 3.5. Let $x_{i_{0}}$ be the expected $t=1$ asset price under the market maker’s prior at $t=0$. For any $W(\,\cdot\,)$, we can find $(a_{i})$ such that
$\!\!\!W(x_{j})$ $=a_{i_{0}}(x_{j}-x_{i_{0}})+\sum_{i<i_{0}}a_{i}(x_{j}-x_{i})_{-}+\sum_{i>i_{0}}a_{i}(x_{j}-x_{i})_{+},\quad\mbox{for $j=1,2,\cdots$}.$ (7)
That is, $W(\,\cdot\,)$ can be replicated by holding $a_{i_{0}}$ shares in the underlying asset, and $a_{i}$ shares in out-of-the-money put (resp. call) options at strikes $x_{i}<x_{i_{0}}$ (resp. $x_{i}>x_{i_{0}}$). Thus, the AD markets can be replaced by an equivalent set of asset and option markets.
The insider learns the true payoff distribution $\eta(\,\cdot\,|s_{k})$ of the underlying asset from his private signal $s_{k}$ and then submits orders for the underlying asset and options. Noise trades follow $\varepsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$ in each market $i$ (for the asset or option at strike $x_{i}$). When the insider submits orders $(a_{i})$, the order flow received by the market maker is given by $a_{i}+\varepsilon_{i}$, for $i=1,2,\cdots.$
The rest of the model proceeds along the same lines as the AD formulation. Conversely, the results obtained in the AD formulation translate to options.
Options in Kyle (1985)
In the options formulation, the single-asset case translates as follows.
Proposition 3.3.
(i) Suppose the payoff distributions specified by the insider’s signal differ only in their expected values, and there are no option markets. In this case, our model reduces to a Kyle (1985) model in which only the underlying asset is traded.
(ii) Conversely, if the expected payoff remains identical across possible payoff distributions, then the insider’s demand for the underlying asset is zero. In other words, exploiting information beyond the expected payoff requires options.
Proposition 3.3(i) applies only when there are no option markets. In other words, the linear equilibrium (5) of Kyle (1985) is not robust with respect to the introduction of options. Options allow the insider to reduce demand optimally for lower-probability strikes, which makes his demand necessarily nonlinear. See Example 7.1 for an illustration.
4 Cross-Market Inference
Intuitively, the market maker infers information from order flow not only within each market but also across markets, as illustrated in Section 2. We now formalize this intuition in the general model.
4.1 The Posterior and Pricing Kernel
Formally, the order flow $\omega$ is an element of the probability space $\Omega$ of countable sequences, endowed with an appropriate probability measure. (For finitely many states, $\Omega$ reduces to a finite-dimensional Euclidean space.) According to the market maker’s belief $\widetilde{W}(\,\cdot\,|\,\cdot\,)$ about how the insider trades, the order flow he receives, conditional on each signal $s_{k}$, is a realization $\omega=(\omega_{i})$ of the stochastic sequence
$\widetilde{W}(x_{i}|s_{k})+\varepsilon_{i},\;i=1,2,\cdots.$ (8)
The market maker applies Bayes’ Rule according to his belief about how the insider trades. His unknown parameter is $s_{k}$, with prior probability $\pi_{0}(s_{k})$, $1\leq k\leq K$. His observed data is order flow $\omega$.
Let $\mathbb{P}_{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}$ and $\mathbb{P}_{\!\scriptscriptstyle 0}$ denote the probability densities specifying the stochastic sequences $(\widetilde{W}(x_{i}|s_{k})+\varepsilon_{i})$ and $(\varepsilon_{i})$, respectively, on $\Omega$. In other words, $\mathbb{P}_{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}(\omega)$ is the likelihood of $\omega$ when the insider’s order is $\widetilde{W}(\,\cdot\,|s_{k})$, while $\mathbb{P}_{\!\scriptscriptstyle 0}(\omega)$ is the likelihood when the insider’s order is zero across markets.
To calculate the conditional likelihood $\mathbb{P}(\omega|s_{k})$ of $\omega$, the market maker takes the likelihood ratio of $\mathbb{P}{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}(\omega)$ over $\mathbb{P}{\scriptscriptstyle 0}(\omega)$,
$\frac{\mathbb{P}_{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}(\omega)}{\mathbb{P}_{\!\scriptscriptstyle 0}(\omega)}=\exp\left(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}-\frac{1}{2}\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}\right).$ (9)
(For finitely many states, this likelihood ratio simplifies to a ratio of normal densities on a finite-dimensional Euclidean space—e.g., see Equation (1) for the three-state case.)
The market maker then applies Bayes’ Rule to obtain his posterior distribution over $s_{k}\in S$, after receiving order flow $\omega$,
[…]
where $C(\omega)$ normalizes this expression to sum to one over $S$. With this posterior, the market maker sets his zero-profit AD prices $P(x_{i}|\omega\kern 0.44998pt;\widetilde{W})$ of (3) for each security $x_{i}$.
Theorem 4.1.
Suppose the market maker holds a belief $\widetilde{W}(\,\cdot\,|\,\cdot\,)$ and receives order flow $\omega$. Then, his posterior distribution over signals is given by
[…]
which is determined by the sufficient statistic $\left(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}\right)_{s_{k}\in S}\in\mathbb{R}^{K}.$
4.2 Sufficient Statistics
Recall the following intuition from Section 2: if the market maker expects the insider to trade the straddle under high volatility, receiving order flow consistent with such a strategy confirms this prior expectation, leading him to conclude that volatility is likely high. More precisely, his sufficient statistic $\Delta(\omega)$, defined in Equation (2), is proportional to the projection coefficient of the order flow $\omega=(\omega_{1},\omega_{2},\omega_{3})$ onto the excess demand he expects from the insider under high vs. low volatility, represented by the vector $(1,0,1)-(0,-1,0)$. A high projection coefficient $\Delta(\omega)$ indicates that the order flow aligns closely with the straddle $(1,0,1)$, leading him to revise his belief towards high volatility.
Theorem 4.1 establishes the generality of this intuition under arbitrary information structure: for each signal $s_{k}$, the market maker projects the order flow $\omega$ onto his prior belief about the insider’s demand conditional on $s_{k}$, $\widetilde{W}(\,\cdot\,|s_{k})$. His posterior probability of each $s_{k}$ is determined by the noise-adjusted projection coefficient
$\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}.$
A high coefficient indicates that the order flow $\omega$ align closely with his belief about the informed demand conditional on $s_{k}$. This leads him to revise his prior probability for $s_{k}$ upwards, and vice versa. Thus, the projection coefficient profile across signals, $\left(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}\right)_{s_{k}\in S}$, serves as his sufficient statistic.
The market maker adjusts for the noise trading intensities $(\sigma_{i})$ in making cross-market inferences. Higher noise levels reduce the informativeness of his sufficient statistic $(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}})_{s_{k}\in S}$. In the limit as $\sigma_{k}\!\rightarrow\!\infty$ for each $k$, this statistic reduces to the trivial zero statistic, and his posterior (11) reduces to the prior $\pi_{0}$, i.e., he infers no information from order flow.
When $X$ is a singleton, Theorem 4.1 reduces to the market maker’s inference in the familiar single-asset case: if the market maker expects the insider to buy when the asset value is high, receiving a buy order would lead him to revise the asset price upwards, and vice versa.
Definition 2.
(i) The noise-adjusted projection coefficient of the market maker’s order flow $\omega$ onto $\widetilde{W}(\,\cdot\,|s_{k})$ is
$\Pi_{mm}(\omega,s_{k};\widetilde{W})=\sum_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}.$ (12)
(ii) The noise-adjusted projection coefficient of an insider portfolio $W(\,\cdot\,)$ onto $\widetilde{W}(\,\cdot\,|s_{k})$ is
$\Pi_{insider}(W,s_{k};\widetilde{W})=\sum_{i}\frac{\widetilde{W}(x_{i}|s_{k})W(x_{i})}{\sigma_{i}^{2}}.$ (13)
Theorem 4.1 established that the projection coefficient profile $\Pi_{mm}(\omega,\,\cdot\,;\widetilde{W})$ is the market maker’s sufficient statistic. We will show in Section 5 that $\Pi_{insider}(W,\,\cdot\,;\widetilde{W})$ serves a similar role for the insider. The difference between the two is that while the market maker only observes the combined order flow $\omega$, the insider knows his portfolio $W$.
$\Pi_{mm}$ and $\Pi_{insider}$ summarize the cross-market dependence between the quantities demanded and prices for each agent, based on their respective information. Using $\Pi_{mm}(\omega,\,\cdot\,;\widetilde{W})$, the market maker compares the order flow $\omega$ with his belief about the insider’s demand, signal by signal, to infer information. Similarly, using $\Pi_{insider}(W,\,\cdot\,;\widetilde{W})$, the insider considers how a portfolio $W$ compares with the market maker’s belief, signal by signal, to evaluate its potential price impact across markets.
5 Cross-Market Price Impact
Given the market maker’s belief $\widetilde{W}(\,\cdot\,|\,\cdot\,)$ about his trading strategy and conditional on signal $s_{k}$, the insider’s portfolio choice problem (4) can be written as
$\max\limits_{{\scriptscriptstyle W(\cdot)}}J(W|s_{k};\widetilde{W})$ $=\max\limits_{{\scriptscriptstyle W(\,\cdot\,)}}\,\underbrace{\left(\sum_{i}W(x_{i})\eta(x_{i}|s_{k})\right)}_{\text{expected payoff}}-\underbrace{\left(\sum_{i}W(x_{i})\overline{P}(x_{i},W;\widetilde{W})\right)}_{\text{expected cost}}.$ (14)
Here, $\overline{P}(x_{i},W;\widetilde{W})\equiv\mathbb{E}^{\mathbb{P}_{\!\scriptscriptstyle W}}[P(x_{i}|\omega\kern 0.44998pt;\widetilde{W})]$ denotes the expected price of security $x_{i}$ under the probability distribution $\mathbb{P}_{\!\scriptscriptstyle W}$ of order flow $\omega$ induced by the insider portfolio $W$. Let $\partial\!\kern 0.80002ptJ(v;W)$ denote the insider’s marginal utility from adding a marginal portfolio $v$ to a portfolio $W$. (In the case of finitely many states, $\partial\!\kern 0.80002ptJ(v;W)$ is the inner product of $v$ and the vector of marginal utilities at $W$131313 In the case of finitely many states, $\partial\!\kern 0.80002ptJ(v;W)$ is the directional derivative of $J(\,\cdot\,|s_{k};\widetilde{W})$ at $W$ in the direction of $v$ on a finite-dimensional Euclidean space.)
5.1 First-Order Condition
In trading to exploit his private information, the insider balances each security’s marginal payoff against its marginal cost. This cost comprises, first, the security’s current price (conditional on no price impact) and, second, its price impact across securities. Theorem 5.1 makes this intuition precise in the general setting.
Theorem 5.1.
(Insider FOC)
The insider’s marginal utility from adding a marginal portfolio $v$ to a portfolio $W$ decomposes into
$\underbrace{\partial\!\kern 0.80002ptJ(v;W)}_{\text{marginal utility}}=\underbrace{\partial\!\kern 0.80002ptJ_{p}(v;W)}_{\text{marginal payoff}}-\underbrace{\big{(}\,\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(v;W)+\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)\,\big{)}}_{\text{marginal cost}}$
where the terms are defined as follows:
$\underbrace{\partial\!\kern 0.80002ptJ_{p}(v;W)}_{\text{marginal payoff}}$ $=\sum_{i}v(x_{i})\eta(x_{i}|s_{k}),$ (15)
$\underbrace{\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(v;W)}_{\begin{subarray}{c}\text{AD term}\end{subarray}}$ $=\sum_{i}v(x_{i})\overline{P}(x_{i},W;\widetilde{W}),$ (16)
$\underbrace{\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)}_{\begin{subarray}{c}\text{price impact term}\end{subarray}}$ $=\sum_{i}W(x_{i})\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}\!\kern-0.20004pt\Bigl{[}\underbrace{{\operatorname{\mathbb{Cov}}}\!\left(\eta(x_{i}|\,\cdot\,),\Pi_{insider}(v,\,\cdot\,;\widetilde{W})\,\bigg{|}\,\omega\right)}_{\text{price impact of $v$ on $x_{i}$ conditional on $\omega$}}\Bigr{]}\kern-0.20004pt.$ (17)
Therefore, any optimal portfolio $W$ must satisfy the first-order condition
$\partial\!\kern 0.80002ptJ_{p}(v;W)=\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(v;W)+\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)\quad\mbox{for all marginal portfolio $v$}.$ (18)
Conditional on signal $s_{k}$, the marginal payoff of a security $x_{i}$ is its true payoff $\eta(x_{i}|s_{k})$ known to the insider. Indeed, Equation (15) shows that the marginal payoff functional $\partial\!\kern 0.80002ptJ_{p}(\,\cdot\,;W)$ is identified with the true payoff distribution $\eta(\,\cdot\,|s_{k})$. The first-order condition (18) equalizes this marginal payoff $\partial\!\kern 0.80002ptJ_{p}(v;W)$ with the marginal cost consisting of two terms:
  • AD Term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(v;W)$: This term is the cost of adding the marginal portfolio $v$ assuming no price impact. Indeed, Equation (16) shows that $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(\,\cdot\,;W)$ is identified with the current AD market prices $\overline{P}(\,\cdot\,,W;\widetilde{W})$ set by the market maker.
  • Price Impact Term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$: This term is the cost incurred due to the market maker adjusting prices across markets in response to the insider adding marginal portfolio $v$.
5.2 The Price Impact Term
From the expression of $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ in (17), it follows that the price impact of the portfolio $v$ on security $x_{i}$, conditional on $\omega$, is
${\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\Pi_{insider}(v,\,\cdot\,;\widetilde{W})\,\bigg{|}\,\omega\right).$ (19)
This is the covariance between the payoff profile $\eta(x_{i}|\,\cdot\,)$ and the projection coefficient profile $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})$ across signals, under the market maker’s posterior. The expected price impact of $v$ on $x_{i}$ is given by the expected value of (19) under $\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[\,\cdot\,]$.
When the portfolio $v$ consists of one share of security $x_{j}$, $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})$ is equal to $\frac{\widetilde{W}(x_{j}|\,\cdot\,)}{\sigma_{j}^{2}}.$ Substituting this into (19) yields the following characterization of the cross price impact of $x_{j}$ on $x_{i}$.
Corollary 5.2.
Given market maker belief $\,\widetilde{W}$, the cross price impact of $x_{j}$ on $x_{i}$, conditional on $\omega$, is
$\frac{\partial}{\partial W(x_{j})}P(x_{i},\omega\kern 0.44998pt;\widetilde{W})=\frac{1}{\sigma_{j}^{2}}\cdot{\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\widetilde{W}(x_{j}|\,\cdot\,)\,\bigg{|}\,\omega\right).$ (20)
The expected cross price impact is given by
$\frac{\partial}{\partial W(y)}\overline{P}(x,W;\widetilde{W})=\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}\!\kern-0.80002pt\Bigl{[}\frac{\partial}{\partial W(y)}P(x,\omega\kern 0.44998pt;\widetilde{W})\Bigr{]}.$ (21)
The intuition underlying the cross price impact characterization in Equation (20) is straightforward: demand for $x_{j}$ increases the price of $x_{i}$ if the market maker believes that a high demand for $x_{j}$ correlates with a high payoff of $x_{i}$—in other words, if he believes there is positive correlation between the insider’s demand $\widetilde{W}(x_{j}|\,\cdot\,)$ for $x_{j}$ and the payoff $\eta(x_{i}|\,\cdot\,)$ of $x_{i}$ across signals, and vice versa. A greater noise trading intensity $\sigma_{j}$ for $x_{j}$ reduces this price impact.
Example 5.3.
Suppose a security $x_{j}$ pays off only under signal $s_{k_{j}}$ and not any other signal. Suppose further that the market maker reasonably believes that the insider’s demand for $x_{j}$ is positive under $s_{k_{j}}$ but zero otherwise. Consider the following two scenarios for another security $x_{i}$:
  1. (i)
    (Positive Price Impact) Suppose $x_{i}$ also pays off only under $s_{k_{j}}$, the same signal that triggers payoff for $x_{j}$. Then $x_{j}$ has a positive price impact on $x_{i}$: a strong demand for $x_{j}$ suggests to the market maker that $x_{i}$ is also likely to pay off, leading him to raise the price of $x_{i}$. Indeed, here the payoff $\eta(x_{i}|\,\cdot\,)$ and $\widetilde{W}(x_{j}|\,\cdot\,)$ are positively correlated, both being positive at $s_{k_{j}}$ and zero otherwise. Therefore, by Corollary 5.2, $\frac{\partial}{\partial W(x_{j})}P(x_{i},\omega\kern 0.44998pt;\widetilde{W})>0$.
  2. (ii)
    (Negative Price Impact) Suppose $x_{i}$ only pays off under a signal different from $s_{k_{j}}$. By Corollary 5.2, $x_{j}$ has a negative price impact on $x_{i}$. The intuition is the flip side of scenario (i).
5.3 No-Arbitrage
No-arbitrage is a general necessary condition for the existence of equilibrium. We now state this condition for our setting. Here, once again, the projection coefficient profile $\Pi_{insider}$ between the insider’s trades and the market maker’s belief naturally occurs. Given a market maker belief $\widetilde{W}$, we define the corresponding zero price impact portfolios as portfolios $W$ for which $\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0$. Intuitively, such a portfolio causes zero price impact because it is orthogonal to the market maker’s belief about the insider’s trading strategy. Such a portfolio must yield a zero payoff for the insider; otherwise, arbitrage opportunities arise for the insider. This is Theorem 5.4.
Theorem 5.4.
(No-Arbitrage)
Given any market maker belief, an optimal portfolio for the insider can exist only if the zero price impact portfolios corresponding to that belief yield zero payoff for the insider. Otherwise, the insider can obtain unbounded utility.
6 Equilibrium
Assumption 1.
(w.l.o.g.)   The market maker has a uniform prior on $S$.
Assumption 1 is made without loss of generality. Any other prior $\pi_{0}$ can be reinterpreted as if it were the equivalent case of the uniform prior $\pi_{u}$ and modified security payoffs $\{\frac{\pi_{0}(s_{k})}{\pi_{u}(s_{k})}\eta(\,\cdot\,|s_{k})\}_{s_{k}\in S}$. [^14] Our proofs go through for any payoff function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the modified payoffs. Once equilibrium is established under the uniform prior, this equivalence yields the comparative statics for varying the prior. For instance, if two securities both yield higher payoffs under some $s_{k}$ compared to the other signals, then their cross price impact increases with the prior probability of $s_{k}$. This generalizes from the simplified setting in Section 2 (see footnote 8).
6.1 Canonical Game
In equilibrium, the insider’s zero payoff portfolios must coincide with the zero price impact portfolios. Zero payoff implies zero price impact because the market maker’s equilibrium belief is correct per Definition 1. Conversely, by the no-arbitrage condition of Theorem 5.4, zero price impact implies zero payoff.
The zero payoff portfolios are those orthogonal to the linear span of $\{\eta(\,\cdot\,,s_{k})\}_{k}$—any portfolio $W$ satisfying $\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k})=0$ for all $s_{k}$ yields zero payoff for the insider. Thus, the insider’s payoff-relevant portfolios lie within the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$. We state this result as Proposition 6.1(i). On the other hand, the market maker’s noise-adjusted equilibrium beliefs must contain this span; otherwise, zero price impact allows arbitrage by the insider. This is Proposition 6.1(ii).
Proposition 6.1.
(i) The insider’s equilibrium demand $W^{*}(\,\cdot\,|s_{k})$ conditional on each $s_{k}$ can be restricted to the linear span of $\{\eta(\,\cdot\,|s_{l})\}_{l}$.
(ii) Conversely, each $\eta(\,\cdot\,|s_{k})$ must lie in the linear span of $\{\frac{W^{*}(\,\cdot\,|s_{l})}{\sigma_{\,\cdot\,}^{2}}\}_{l}$ to prevent arbitrage opportunities.
To make the reduction of Proposition 6.1 explicit, represent a portfolio $W(\,\cdot\,)=\sum\limits_{k=1}^{K}d_{k}\eta(\,\cdot\,|s_{k})$ in the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$ by the vector $d=(d_{k})\in\mathbb{R}^{K}$ of its expansion coefficients. Similarly, represent a market maker belief $\widetilde{W}(\,\cdot\,|\,\cdot\,)$ in the same linear span,
$\widetilde{W}(\,\cdot\,|s_{k})=\sum_{l=1}^{K}\tilde{d}^{(k)}_{l}\eta(\,\cdot\,|s_{l}),\;\tilde{d}^{(k)}\in\mathbb{R}^{K}\;\mbox{ for }1\leq k\leq K,$
by the matrix $\widetilde{D}=[\tilde{d}^{(1)}\;\cdots\;\tilde{d}^{(K)}]\in\mathbb{R}^{K\times K}$.
Definition 3.
Define the information intensity matrix $\mathbf{L}\in\mathbb{R}^{K\times K}$ of the trading game by [^15] In other words, ${\bf L}$ is the positive-semidefinite square root of $\left[\sum\limits_{i}\frac{\eta(x_{i}|s_{k})\eta(x_{i}|s_{l})}{\sigma_{i}^{2}}\right]_{1\leq k,l\leq K}$. It is positive definite when the $\eta(\,\cdot\,|s_{k})$’s are linearly independent.
${\bf L}^{2}=\left[\sum_{i}\frac{\eta(x_{i}|s_{k})\eta(x_{i}|s_{l})}{\sigma_{i}^{2}}\right]_{1\leq k,l\leq K}\;\;\mbox{ and }\;\;{\bf L}^{T}={\bf L}.$
${\bf L}$ generalizes the noise-adjusted variation $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ from the single-asset setting (6). A larger $\bf{L}$ means greater noise-adjusted (co-)variations in payoffs across securities and signals, reflecting higher information intensity. [^16] Here, ${\bf L}$ is ordered based on positive-semidefiniteness. This is a partial order. Unlike the single-asset setting, the information intensity across securities and signals is not always directly comparable between two specifications.
By Proposition 6.1, the columns of ${\bf L}$ span the expansion coefficients of the insider’s payoff-relevant portfolios in equilibrium. Therefore, we can apply the canonical transformation
$d\mapsto\hat{d}={\bf L}d.$ (22)
Under this transformation, a market maker belief $\widetilde{D}$ transforms as
$\widehat{D}={\bf L}\widetilde{D}=[\hat{d}^{(1)}\cdots\hat{d}^{(K)}].$ (23)
Canonical Transformation: Intuition
The canonical transformation (22) normalizes the trading game by the insider’s information intensity. For intuition, suppose the $\eta(\,\cdot\,|s_{k})$’s have non-overlapping payoffs across states, such as in Section 2. In this case, there is zero covariation of payoffs across signals, which results in a diagonal information intensity matrix ${\bf L}$. The $k$-th diagonal entry of ${\bf L}$ is $c_{k}^{2}=\sum\limits_{i}\frac{\eta(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}$, which is the noise-adjusted payoff variation across securities conditional on $s_{k}$. The canonical transformation rescales each $\eta(\cdot|s_{k})$ by $\frac{1}{c_{k}}$, i.e., it normalizes the payoff variations to one across signals, thereby transforming ${\bf L}$ to the identity matrix. [^17] We note again that our proofs apply to any positive function $\eta(\,\cdot\,|\,\cdot\,)\colon X\times S\rightarrow\mathbb{R}$, including the rescaled payoffs. This normalization yields an isomorphic game that is invariant with respect to $\eta(\,\cdot\,|\,\cdot\,)$ and $(\sigma_{i})$ and symmetric across signals.
Theorem 6.2.
Under Assumption 1, the Bayesian trading game between the insider and market maker is isomorphic to the canonical game, a pseudo-trading game defined as follows:
  • There are markets for pseudo-securities $k=1,\cdots,K$.
  • Conditional on signal $s_{k}$, pseudo-security $k$ has payoff of $1$ while all others have payoffs of 00.
  • The insider observes the signal, while the market maker has a uniform prior on signals.
  • The insider submits orders $\hat{d}\in\mathbb{R}^{K}$ for the pseudo-securities, and the market maker receives order flow $\hat{\omega}=\hat{d}+\widehat{N}$, where $\widehat{N}_{k}$ are i.i.d. $\mathcal{N}(0,1)$ noise trades for $k=1,\cdots,K$.
  • The market maker has a belief $\widehat{D}$ as specified in (23). Upon receiving $\hat{\omega}$, he updates the posterior probability of each signal $s_{k}$ (which is also the pseudo-security $k$ price) to
    $\hat{\pi}_{1}(k|\hat{\omega}\kern 0.44998pt;\widehat{D})\propto e^{(\hat{d}^{(k)})^{T}\hat{\omega}-\frac{1}{2}(\hat{d}^{(k)})^{T}\hat{d}^{(k)}}.$ (24)
  • Conditional on $s_{k}$ and given market maker belief $\widehat{D}$, the insider maximizes expected profit:
    $\max_{\hat{d}\,\in\,\mathbb{R}^{K}}\,\hat{d}_{k}-\hat{\pi}_{1}(\hat{d}\kern 0.44998pt;\widehat{D})^{T}\hat{d}\equiv\max_{\hat{d}\,\in\,\mathbb{R}^{K}}J(\hat{d}|k\kern 0.44998pt;\widehat{D})$ (25)
    where $\hat{\pi}_{1}(\hat{d}\kern 0.44998pt;\widehat{D})$ is the expected value of the prices (24) over the distribution of $\hat{\omega}$.
Equilibrium in Canonical Game
In the canonical game, an equilibrium is specified by a $K\times K$ strategy matrix $D^{*}=[\delta^{(1)}\;\cdots\;\delta^{(K)}]$ such that, under the same market maker belief $D^{*}$, each strategy $\delta^{(k)}\in\mathbb{R}^{K}$ solves the insider’s problem (25) conditional on $s_{k}$, i.e.,
$\delta^{(k)}\in\operatorname*{argmax}_{\hat{d}\,\in\,\mathbb{R}^{K}}J(\hat{d}|k\kern 0.44998pt;D^{*}),\;\mbox{for each}\;k=1,\cdots,K.$ (26)
Corollary 6.3.
Given the isomorphism between the two games, an equilibrium $D^{*}$ of the canonical game maps to the equilibrium of the original trading game (see Definition 1) specified by
$W^{*}(\,\cdot\,|s_{k})=\sum_{l=1}^{K}\beta^{(k)}_{l}\eta(\,\cdot\,|s_{l}),\,\,\mbox{ where each $\beta^{(k)}$ is the $k$-th column of the matrix ${\bf L}^{-1}D^{*}$.}$ (27)
We refer to the matrix ${\bf L}^{-1}D^{*}$ as the canonical form of the original game equilibrium (27).
Remark 6.4.
When the $\eta(\,\cdot\,|s_{k})$’s are linearly dependent, ${\bf L}^{-1}$ is interpreted as the pseudoinverse of ${\bf L}$. Compared to the case of linearly independent $\eta(\,\cdot\,|s_{k})$’s, the only caveat is that the informed demand portfolio $W^{*}(\,\cdot\,|s_{k})$ of (27) is unique only up to its payoff, but not the allocations of its constituent $\eta(\,\cdot\,|s_{l})$’s.
Insider FOC in Canonical Game
Let $p\in[0,1]^{I}$ denote the random vector (24) of the market maker’s posterior in the canonical game (with the dependence on $\hat{\omega}$ and $\widehat{D}$ understood) and $\mbox{diag}(p)\in\mathbb{R}^{K\times K}$ denote the matrix with $p$ along the diagonal and zeros elsewhere. Then, conditional on observing $s_{k}$ and given market maker belief $\widehat{D}$, the first-order condition for the insider’s problem (25) in the canonical game is
[…]
where $e_{k}$ denotes the $k$-th standard basis vector, and the expectation $\mathbb{E}^{(\hat{d};\widehat{D})}[\,\cdot\,]$ is taken over the distribution of $\hat{\omega}$ induced by the insider’s choice $\hat{d}$ when the market maker holds belief $\widehat{D}$. The terms in the first-order condition (28) are the isomorphic counterparts to those in Theorem 5.1.
6.2 Symmetric Equilibrium
The canonical game is symmetric—up to permutation on $k$, the game is identical conditional on each signal $s_{k}$. This symmetry suggests an equilibrium structure of the form
$D^{*}=\alpha{\bf Q},\mbox{ for some }\alpha>0$ (29)
where ${\bf Q}=\mathbf{I}-\frac{1}{K}\bar{e}\bar{e}^{T}\in\mathbb{R}^{K\times K}$, with $\mathbf{I}$ denoting the identity matrix and $\bar{e}$ the vector of $1$’s.
This means that, conditional on $s_{k}$, the insider buys $\alpha(1-\frac{1}{K})$ shares of pseudo-security $k$ and sells the other pseudo-securities for $\frac{\alpha}{K}$ shares each. This demand corresponds to the $k$-th column of $D^{*}$ in (29). By substituting this equilibrium postulate into the first-order condition (28) conditional on each signal $s_{k}$ and collating the resulting vector equations side-by-side, we derive a matrix equation
$\Phi(\alpha){\bf Q}=0$ (30)
for some scalar function $\Phi\colon[0,\infty)\rightarrow\mathbb{R}$. We prove that each first-order condition is sufficient for optimality at such a fixed point of the collated first-order conditions (30). Thus, an equilibrium corresponds to a solution $\alpha^{*}>0$ to the equation $\Phi(\alpha)=0$. This is Theorem 6.5.
Theorem 6.5.
There exists $\alpha^{*}>0$ such that $\alpha^{*}{\bf Q}$ is an equilibrium of the canonical game. Equivalently, $\alpha^{*}{\bf L}^{-1}{\bf Q}$ specifies an equilibrium of the original trading game, in the canonical form of Corollary 6.3.
Example 6.6.
(Theorem 6.5, Binary Signal Case)
Suppose the signal is binary, $S=\{s_{1},s_{2}\}$. Substituting the equilibrium postulate (29) $D^{*}=\frac{1}{2}\begin{bmatrix}\alpha&-\alpha\\ -\alpha&\alpha\end{bmatrix}$ into the insider’s first-order condition (28) conditional on, say, $s_{1}$ gives
$\begin{bmatrix}1\\ 0\end{bmatrix}-\begin{bmatrix}\mathbb{E}[p_{1}]\\ 1-\mathbb{E}[p_{1}]\end{bmatrix}-\alpha^{2}\begin{bmatrix}\mathbb{E}[p_{1}p_{2}]\\ -\mathbb{E}[p_{1}p_{2}]\end{bmatrix}=(1-\mathbb{E}[p_{1}]-\alpha^{2}\mathbb{E}[p_{1}p_{2}])\begin{bmatrix}1\\ -1\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$ (31)
where $\mathbb{E}[\,\cdot\,]$ is the expectation taken under $D^{*}$, and $p_{k}$ is the posterior probability of $s_{k}$, $k=1,2$. Substituting $D^{*}$ into the market maker’s posterior (24) gives
$(p_{1},p_{2})\stackrel{{\scriptstyle d}}{{\sim}}(\frac{e^{Z}}{e^{Z}+1},\frac{1}{e^{Z}+1}),\quad\mbox{ where }Z\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(\frac{\alpha^{2}}{2},2\alpha^{2}).$ (32)
By (32), the quantities $\mathbb{E}[p_{1}]$ and $\mathbb{E}[p_{1}p_{2}]$ in (31) are moments of a logit-normal distribution and functions of $\alpha$. Write $\mathbb{E}[p_{1}]$ as $\phi_{1}(\alpha)$ and $\mathbb{E}[p_{1}p_{2}]$ as $\phi_{2}(\alpha)$. Define $\Phi(\alpha)=1-\phi_{1}(\alpha)-\alpha^{2}\phi_{2}(\alpha)$. Then the first-order conditions (31) conditional on $s_{1}$ and $s_{2}$ become, respectively,
$\Phi(\alpha)\begin{bmatrix}1\\ -1\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}\,\,\mbox{ and }\,\,\Phi(\alpha)\begin{bmatrix}-1\\ 1\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}.$
[…]
Intuition of $\alpha^{*}$
The endogenous constant $\alpha^{*}$ characterizing equilibrium is the insider’s optimal portfolio size in the canonical pseudo-trading game. The intuition behind $\alpha^{*}$ builds on the simplified setting discussed in Section 2—see again Figure 1b. This is demonstrated by Example 6.7.
Example 6.7.
In the same setting as Example 6.6, consider the insider’s utility and the market maker’s posterior (32), as $\alpha$ varies.
  • When $\alpha=0$, the insider has zero demand, yielding zero utility and revealing no information. The market maker’s posterior reduces to the prior. This cannot be an equilibrium—the insider’s marginal utility is strictly positive.
  • When $\alpha\rightarrow\infty$, the insider scales up his demand without bound, fully revealing his information and resulting in zero profit. The market maker’s posterior concentrates on the insider’s signal. This cannot be an equilibrium—the insider’s marginal utility is strictly negative.
  • When $\alpha=\alpha^{*}$, the insider’s marginal utility is zero, the market maker’s belief is correct, and equilibrium is obtained. The equilibrium posterior is biased towards the insider’s signal, partially revealing his information because the market maker correctly anticipates his trades.
Comparative Statics of $\alpha^{*}$
[…]
[5pt] 02040608010000.511.52 $x$-axis: number of signals, $y$-axis: $\alpha^{*}$
Figure 2: Comparative Statics of Endogenous Constant $\alpha^{*}$
$\alpha^{*}$ is determined only by the number of signals $K$—it is independent of $\eta(\,\cdot\,|\,\cdot\,)$ or $(\sigma_{i})$. $\alpha^{*}(K)$ is an increasing concave function of $K$, as shown in Figure 2. Increasing the number of signals enhances the insider’s informational advantage over the market maker, resulting in larger insider trades. However, larger insider trades also introduce greater variation in informed demand across signals, diminishing the incremental gain in the insider’s information advantage.
7 Price Discovery
7.1 Informed Demand
The equilibrium informed demand $W^{*}(\cdot|s_{k})$, conditional on signal $s_{k}$, is
$W^{*}(\,\cdot\,|s_{k})=\sum_{l=1}^{K}\beta^{(k)}_{l}\eta(\,\cdot\,|s_{l}),\;\;\mbox{where $\beta^{(k)}$, the $k$-th column of $\alpha^{*}{\bf L}^{-1}{\bf Q}$, defines the allocation weights.}$. (33)
Applying the replication formula (7) yields the equivalent options portfolio.
Informed Demand Portfolio Construction
The informed demand portfolio (33) is constructed following these simple steps:
  • Step 1: Form the initial portfolio by buying $1-\frac{1}{K}$ shares of the observed distribution $\eta(\,\cdot\,|s_{l})$ and selling $\frac{1}{K}$ shares of each remaining distribution $\eta(\,\cdot\,|s_{l})$, $l\neq k$.
  • Step 2: Adjust the initial portfolio allocations by applying the linear transformation ${\bf L}^{-1}$.
  • Step 3: Scale the portfolio by the endogenous constant $\alpha^{*}$.
Step 1 forms the initial portfolio according to the allocation weights specified by the $k$-th column of ${\bf Q}$. Step 2 adjusts the allocations to account for the insider’s information intensity ${\bf L}$; higher information intensity forces the insider to reduce the size of his trades across securities and signals. Step 3 scales the portfolio by $\alpha^{*}$ to optimize with respect to the market maker’s prior uncertainty over signals (see again Figure 2).
Single-Asset vs. AD Informed Demand
The AD informed demand (33), in its canonical form as per Corollary 6.3, shares a similar intuition with the single-asset informed demand (5):
informed $\text{{\em informed demand}}\,\begin{cases}\text{{\em single-asset} :}&\frac{\sigma_{\varepsilon}}{\sigma_{v}}\cdot(s-v_{0})\\[2.58334pt] \text{{\em contingent claims} :}&{\bf L}^{-1}\cdot\alpha^{*}{\bf Q}\end{cases}.$
  • Information Intensity ($\frac{\sigma_{\varepsilon}}{\sigma_{v}}$ vs. ${\bf L}$): Both $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ and ${\bf L}$ quantify the insider’s information intensity in their respective settings. In the single-asset setting, a larger $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ means the market maker can more effectively infer the variation in the asset payoff from the variation in order flow, thus reducing informed demand. For contingent claims, a larger ${\bf L}$ means the market maker can better infer the variations in the payoff distributions across signals from the variations in cross-market order flow, which reduces informed demand.
  • Signal Influence ($(s-v_{0})$ vs. $\alpha^{*}{\bf Q}$): In the single-asset setting, $(s-v_{0})$ scales an initial one-share portfolio based on the signal’s deviation from the market maker’s prior mean. For contingent claims, $\alpha^{*}$ scales the initial long-short portfolio of ${\bf Q}$ according to the market maker’s prior uncertainty regarding the signal.
The following examples illustrate the informed demand for trading on the first, second, and third moments of an underlying asset payoff. These examples explain how the corresponding option strategies used in trading practice emerge as equilibrium strategies. Other common option strategies are explained similarly under appropriate specifications. In Examples 7.1, 7.2, and 7.3, the signal is binary, i.e., $S=\{s_{1},s_{2}\}$, and all payoff distributions are suitably discretized.
(a) High expected payoff signal $s_{1}$
($x$-axis: state/option strike, $y$-axis: probability)
(b) Informed demand $W^{*}(\,\cdot\,|s_{1})$—Bull Spread
($x$-axis: state/option strike, $y$-axis: payoff)
(c) Low expected payoff signal $s_{2}$
(d) Informed demand $W^{*}(\,\cdot\,|s_{2})$—Bear Spread
Figure 3:
Bull/Bear Spreads The left column illustrates the insider’s private signal, indicated by the shaded distributions. The right column depicts the informed demand $W^{*}$ conditional on the respective signals.
(a)/(b): high expected payoff signal/bull spread, (c)/(d): low expected payoff signal/bear spread.
Example 7.1.
(Trading on Mean - Bull/Bear Spread)
Let signals $s_{1}$ and $s_{2}$ specify normal payoff distributions with respective means $\mu_{1}>\mu_{2}$ and an identical variance. Observing $s_{1}$ (resp. $s_{2}$) informs the insider of a high (resp. low) expected payoff. This is a Kyle (1985)-type setting, where private information concerns the expected payoff, but the insider here can trade derivatives.
The left column of Figure 3 shows the possible payoff distributions, with the insider’s private signal shaded. The right column of Figure 3 shows the insider’s demand corresponding to each signal. These nonlinear informed demands require options. Indeed, they correspond to well-known option strategies (see Hull (2015)). When the insider anticipates high expected payoff, his option strategy in Figure 3b takes the form of a bull spread. When he anticipates low expected payoff, his option strategy in Figure 3d takes the form of a bear spread. These option strategies maximize gains in the anticipated direction (up or down) while limiting exposure at extreme out-of-the-money strikes that are unlikely to pay off. In contrast, the linear demand in Kyle (1985) is sub-optimal because it entails unnecessarily large bets at the extreme strikes.
(a) High volatility signal $s_{1}$
(b) Informed demand $W^{*}(\,\cdot\,|s_{1})$—Straddle
(c) Low volatility signal $s_{2}$
(d) Informed demand $W^{*}(\,\cdot\,|s_{2})$—Butterfly
Figure 4:
Straddle/Butterfly The left column illustrates the insider’s private signal, indicated by the shaded distributions. The right column depicts the informed demand $W^{*}$ conditional on the respective signals.
(a)/(b): high volatility signal/straddle, (c)/(d): low volatility signal/butterfly.
Price Impact Three states are indicated—$x_{1}$, $x_{2}$, and $x_{3}$.
$\bullet\,x_{2}$: zero price impact on all securities because its payoff is constant (has zero variation) across signals. Thus, the informed demand for $x_{2}$ must be zero—see (b) and (d)
$\bullet\,x_{1}$ and $x_{3}$: positive cross price impact because they have positively correlated (in fact, identical) payoffs.
Example 7.2.
(Trading on Volatility - Straddle/Butterfly)
Let signals $s_{1}$ and $s_{2}$ specify normal payoff distributions with respective variances $\sigma_{1}^{2}>\sigma_{2}^{2}$ and an identical mean. Observing $s_{1}$ (resp. $s_{2}$) informs the insider of high (resp. low) volatility.
The left column of Figure 4 shows the possible payoff distributions, with the insider’s private signal shaded. The right column of Figure 4 shows the informed demand corresponding to each signal. The informed demands align with common option strategies for volatility trading (see Hull (2015)). When the insider anticipates high volatility, his option strategy in Figure 4b resembles a straddle or long iron condor pattern. When he anticipates low volatility, his option strategy in Figure 4d follows a butterfly pattern.
(a) Right-skewed signal $s_{1}$
(b) Informed demand $W^{*}(\,\cdot\,|s_{1})$
(c) Left-skewed signal $s_{2}$
(d) Informed demand $W^{*}(\,\cdot\,|s_{2})$
Figure 5: Trading on Skewness The left column illustrates the insider’s private signal, indicated by the shaded distributions. The right column depicts the informed demand $W^{*}$ conditional on the respective signals.
Example 7.3.
(Trading on Skewness)
Let signals $s_{1}$ and $s_{2}$ specify skew-normal payoff distributions with respective skewness parameters $\alpha_{1}>\alpha_{2}$. All moments other than skewness are identical across signals. The left column of Figure 5 shows the possible payoff distributions, with the insider’s private signal shaded. The right column of Figure 5 shows the informed demand corresponding to each signal. The informed demand is implemented by long-short option strategies, with the portfolio’s long position aligned with the skewness direction. [^19] In practitioner parlance, these option strategies are called “Christmas tree spreads” (see Vine (2011)).
7.2 Price Impact
In Corollary 5.2, substituting $W^{*}$ for $W$ and $\widetilde{W}$ in (21) yields the equilibrium expected price impact $\Lambda_{i,j}$ between securities $x_{i}$ and $x_{j}$:
$\Lambda_{i,j}=(1-\frac{1}{K})\cdot\alpha^{*}\cdot\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}}\cdot\mathbb{E}\left[{\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\eta(x_{j}|\,\cdot\,)\big{|}\omega\right)\right]$ (34)
where ${\operatorname{\mathbb{Cov}}}(\,\cdots\,,\,\cdots\,\big{|}\omega)$ is the covariance over signals $s_{k}\in S$ under the market maker’s equilibrium posterior, conditional on order flow $\omega$. The expectation $\mathbb{E}[\,\cdot\,]$ is taken over $\omega$ with respect to its equilibrium distribution. By removing this expectation in (34), we obtain the equilibrium price impact conditional on order flow $\omega$.
Complementary Factors of Price Impact
The equilibrium cross price impact $\Lambda_{i,j}$ between securities $x_{i}$ and $x_{j}$ is determined by their payoff covariance, adjusted for factors related to the informed and noise trades:
  1. (i)
    Portfolio Allocation $(1-\frac{1}{K})$: The weight of the true distribution in the informed demand (33).
  2. (ii)
    Informed Demand Scale $(\alpha^{*}$): As the informed demand (33) scales up by $\alpha^{*}$, price impact responds one-to-one.
  3. (iii)
    Noise Intensity Adjustment $(\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}})$: Higher noise trading intensity in either market reduces cross price impact.
  4. (iv)
    Payoff Covariance $(\mathbb{E}\left[{\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\eta(x_{j}|\,\cdot\,)\big{|}\omega\right)\right])$: The (expected) covariance of securities $x_{i}$ and $x_{j}$ payoffs under the market maker’s equilibrium posterior.
Thus, $\Lambda_{i,j}$ extends the intuition behind cross-market price impact discussed in Section 2.
Example 7.4.
  1. (i)
    (Negative Cross Price Impact.) If security $x_{i}$ pays off only under signal $s_{k_{i}}$ and $x_{j}$ only under $s_{k_{j}}\neq s_{k_{i}}$, their equilibrium cross price impact is negative. This is the same scenario as Example 5.3(ii), but in equilibrium. Upon receiving a buy order for $x_{i}$, the market maker infers that $x_{j}$ is unlikely to pay off and lowers its price accordingly, and vice versa. Indeed,
    $\Lambda_{i,j}=-(1-\frac{1}{K})\cdot\alpha^{*}\cdot\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}}\cdot\mathbb{E}[\eta(x_{i}|s_{k_{i}})\pi^{*}_{1}(s_{k_{i}}|\omega)\cdot\eta(x_{j}|s_{k_{j}})\pi^{*}_{1}(s_{k_{j}}|\omega)]<0.$
  2. (ii)
    (Zero Cross Price Impact.) If a security $x_{j}$’s payoff has zero variation across signals (i.e., $\eta(x_{j}|\,\cdot\,)$ is constant), then its order flow has zero price impact across markets. In turn, no-arbitrage implies that informed demand for $x_{j}$ must be zero. Conversely, if the informed demand for $x_{j}$ is zero across signals, its order flow is pure noise and, therefore, has zero price impact across markets. This is illustrated by the security $x_{2}$ shown in Figure 4c. Figures 4b and 4d confirm the corresponding informed demand $W^{*}(x_{2}|\,\cdot\,)=0$ to be indeed zero across signals.
  3. (iii)
    (Positive Cross Price Impact.) If two securities have identical payoffs, then their cross price impact is naturally positive. This is illustrated by the securities $x_{1}$ and $x_{3}$ shown in Figure 4c.
We obtain the following general formula for cross price impact between derivatives.
Corollary 7.5.
(Price Impact Between Derivatives)
Let $\varphi_{1},\varphi_{2}\colon X\rightarrow\mathbb{R}$ represent state-contingent claims (e.g. options). Then their cross price impact is
$\sum_{i,j}\varphi_{1}(x_{i})\varphi_{2}(x_{j})\Lambda_{i,j}$ (35)
Applying Corollary 7.5 to Example 7.2 yields the following result.
Proposition 7.6.
Under the specification of Example 7.2, the cross price impact between the put-call pair in a straddle is higher conditional on the high-volatility signal than on the low-volatility signal.
In other words, a high cross price impact between the put-call pair in a long-volatility straddle means that volatility is underpriced and should therefore correct upwards as volatility price discovery takes place. These insights regarding cross-market information dynamics generalize to higher moments beyond volatility, offering testable empirical predictions in a broad framework (see Section 8).
120.00.40.8
(a) $K=2$
12340.00.40.8
(b) $K=4$
1234560.00.40.8
(c) $K=6$
123456780.00.40.8
(d) $K=8$
Figure 6: Comparative Statics - Market Maker’s Posterior ($x$-axis: $S=\{s_{1},\cdots,s_{K}\}$, $y$-axis: probability) These graphs illustrate the market maker’s expected posterior probabilities on the signal space $S$ conditional on the insider observing signal $s_{1}$.
7.3 Information Efficiency of Prices
The equilibrium posterior distribution of the market maker over order flow $\omega$ is characterized as follows.
Proposition 7.7.
When the insider observes $s_{k}$, the market maker’s equilibrium posterior $\pi^{*}_{1}(ds,\omega)$ follows a logistic-normal distribution
$q^{(k)}\propto(e^{Z_{1}},\cdots,e^{(\alpha^{*})^{2}+Z_{k}},\cdots,e^{Z_{K}})^{T},\;\;\mbox{ where }\;(Z_{k})_{k=1}^{K}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,(\alpha^{*})^{2}\,{\bf Q}).$ (36)
The equilibrium distribution of AD prices is an immediate corollary (the equilibrium $\omega$-by-$\omega$ pricing kernel can be characterized similarly):
Corollary 7.8.
When the insider observes $s_{k}$, the equilibrium AD prices $P^{*}(\,\cdot\,|\omega)$ are distributed as $P^{*}(\,\cdot\,|\omega)\stackrel{{\scriptstyle d}}{{\sim}}\sum\limits_{l=1}^{K}q_{l}^{(k)}\eta(\,\cdot\,|s_{l})$, with $q^{(k)}$ as defined in Proposition 7.7. Therefore, the expected equilibrium AD prices, $\overline{P}^{*}(\,\cdot\,)$, are given by
$\overline{P}^{*}(\,\cdot\,)=\sum_{l=1}^{K}\mathbb{E}[q_{l}^{(k)}]\eta(\,\cdot\,|s_{l}).$ (37)
Thus, the information efficiency of AD prices is characterized by the weight, $\mathbb{E}[q_{k}^{(k)}]$, of the true payoff distribution $\eta(\,\cdot\,|s_{k})$ in the equilibrium prices (37).
Remark 7.9.
In our setting, information efficiency encompasse higher-moment information across multiple securities, unlike Kyle (1985), which considers only the first moment and a single asset. In Kyle (1985), information efficiency is measured by the prior-to-posterior variance ratio of the asset price, where a higher ratio indicates that a larger fraction of the insider’s information is reflected in the price. This single-asset measure is inadequate for our setting. Security-by-security variance ratios do not reflect how much higher-moment information is incorporated across security prices. In our setting, information efficiency is directly measured by the fraction of the insider’s information (i.e., true payoffs) in the market prices.
Corollary 7.10.
The information efficiency of AD prices does not depend on the possible distributions across future states $\eta(\,\cdot\,|\,\cdot\,)$ or on the noise trading intensities $(\sigma_{i})$.
Remark 7.11.
Corollary 7.10 points to a basic property of complete markets overlooked in prior literature: the information efficiency of prices is robust to the specifics of information. As the insider’s trades (co-)vary across securities according to his information intensity, the market maker correctly anticipates these (co-)variations and equilibrates prices to incorporate that information. The degree to which this occurs is independent of the specifics of information. This is the price discovery counterpart of efficient risk sharing in complete markets, where prices equilibrate to allocate risk efficiently independent of the initial distribution of risk.
Figure 6 illustrates the expected posterior probabilities $\mathbb{E}[q_{k}^{(1)}]$, $1\!\leq\!k\!\leq\!K$, when the true signal is $s_{1}$, as $K$ increases. It shows that the information efficiency $\mathbb{E}[q_{1}^{(1)}]$ (equal across $k$) is decreasing in the number of signals. More precisely, Figure 7 demonstrates that $\mathbb{E}[q_{k}^{(k)}]$ is a decreasing, convex function of $K$, as stated in Proposition 7.12.
[…]
[5pt] 1204060801000.40.60.81 $x$-axis: number of signals, $y$-axis: information efficiency measure $\mathbb{E}[q_{k}^{(k)}]$
Figure 7: Comparative Statics of Information Efficiency
Proposition 7.12.
A larger number of possible signals leads to lower price information efficiency, with a diminishing marginal effect.
Proposition 7.12 is intuitive and consistent with the comparative statics of $\alpha^{*}$ shown in Figure 2. A larger number of signals makes it more difficult for the market maker to infer information from trades, thereby reducing the information efficiency of prices. The marginal effect on information efficiency is diminishing because the resulting larger trades by the insider, aimed at exploiting the market maker’s prior uncertainty, increase the variation in his trades across signals. This increased variation reveals some of his information, partially offsetting the loss in information efficiency.
8 Empirical Implications
Higher-Moment Adverse Selection
Empirical studies on adverse selection traditionally focus on the within-market price impact, which estimates Kyle’s lambda. [^20] See Glosten and Harris (1988), Lin et al. (1995), Huang and Stoll (1997), Goyenko et al. (2009), Hendershott et al. (2011), and Makarov and Schoar (2020). However, this approach does not capture information about higher moments. For volatility information, the implied volatility from option prices serves as a common proxy for future volatility expectations. [^21] See, for example, Goyal and Saretto (2009). However, implied volatility is based on the assumption of constant volatility and is a result of reduced form model mis-specification. Our characterization of cross price impact extends Kyle’s lambda to an empirical measure that directly reflects higher-moment adverse selection effects across options. This new measure extends existing empirical proxies and can help reduce biases associated with ignoring cross-market effects.
Predictability of Return Higher Moments
The adverse selection measure leads to testable hypotheses about the predictability of higher moments in the underlying asset’s returns. A high cross price impact between the put-call pair in the straddle means that the market maker is adverse-selecting with respect to volatility, indicating that the future volatility of the underlying payoff is currently underpriced. This, in turn, should predict an increase in the volatility of the underlying payoff. This is the empirical content of Proposition 7.6:
Hypothesis 8.1.
Higher cross price impact between the options in a straddle (resp. butterfly) predicts an increase (resp. decrease) in the volatility of the underlying return.
Similarly, under the specification of Example 7.3, a higher cross price impact between the long-short option positions for trading on right-skewness predicts an increase in the skewness of the underlying return. More generally, combining the informed demand formula (33) and the cross price impact formula of Corollary 7.5 yields the following empirical hypothesis.
Hypothesis 8.2.
The cross price impact between options appropriately specified by the informed demand formula predicts the corresponding moments of the underlying return. [^22] This hypothesis may be tested by regressing realized moments (e.g., realized volatility) on the lagged cross price impact between prescribed option contracts (e.g., the put-call pair in a straddle).
If the cross price impact predicts the relevant moment, the corresponding option strategy would generate positive returns, indicating that market mispricings in higher moments are corrected over time. The degree to which this holds empirically can reveal several insights: it might apply only up to a certain moment order, indicating the extent of higher-moment price discovery, or only within specific time horizons, indicating the frequency of such price adjustments. Alternatively, rejection of Hypothesis 8.2 for certain moments would imply that option markets are already efficient with respect to those specific moments.
Question 8.3.
What moments allow the prescribed option strategies in Hypothesis 8.2 to generate positive returns? In other words, to what extent does higher-moment price discovery occur in option markets?
While Conrad et al. (2013) found that option-implied skewness could sometimes be informative for return asymmetries, Hypothesis 8.2 offers a more systematic and granular prediction framework by relating options cross price impact directly with future moments of the underlying return distribution.
Cross-Section of Option Returns
Applying the higher-moment predictability of Hypothesis 8.2 in the cross-sectional, rather than time series, context suggests a new set of candidate higher-moment factors for the cross-section of option returns.
Conjecture 8.4.
Systematic factors for the cross-section of option returns can be constructed the prescribed option portfolios targeting higher moments that yield positive returns. [^23] The candidate factors may be constructed by sorting prescribed option portfolios according to the cross price impact between options in the portfolio and forming long-short portfolios as factors—for example, sorting straddles according to the cross price impact between the constituent put-call pair and forming long-short portfolios.
The cross-section of option returns remains a puzzle—see Bali and Murray (2013), Cao and Han (2013), Christoffersen et al. (2018), and Zhan et al. (2022). This conjecture suggests a new approach that links higher-moment information to the cross-section of option returns. By hypothesizing that long-short portfolios sorted by cross price impact could capture systematic variation in option returns, this set of higher-moment factors is based on information dynamics and distinct from previously considered factors like the underlying stock characteristics or liquidity measures (see Bali and Murray (2013), Cao and Han (2013), and Zhan et al. (2022), and Christoffersen et al. (2018)).
0.010.340.671.001.331.660.20.30.40.50.60.70.8
(a) $x$-axis: log-moneyness, $y$-axis: implied vol
0.010.340.671.001.331.660.20.30.40.50.60.70.8
Figure 8: Insider-Induced Volatility Smile
(a) displays the insider-induced implied volatility across option strikes under Example 7.2.
(b) displays this insider-induced volatility smiles for different values of true volatility: $0.25$ (solid line, same line as (a)), $0.3$ (dashed line), and $0.35$ (dotted line).
Computation: For an order flow realization $\omega$, we use the pricing kernel $P^{*}(,\cdot,|\omega;s_{1})$ to compute the equilibrium option prices across strike $K$. Based on these option prices, we compute the implied volatility $\sigma_{\scriptscriptstyle IV}(\omega,K)$ for each strike $K$. This is repeated for $1,000$ realizations of $\omega$, with the plots displaying the average implied volatilities over log-moneyness. The risk-free rate and dividend yield are set to zero.
Insider-Induced Volatility Smile
Our results show that the volatility smile emerges as volatility information is incorporated into option prices. In Example 7.2, when an insider trades on a high-volatility signal, the implied volatilities of equilibrium option prices produce a smile that replicates the typical observed volatility smile pattern with a skew towards lower strikes (see Figure 8a).
The intuition behind this insider-induced smile is straightforward. Trading on high volatility leads the insider to buy deep in-the-money and out-of-the-money options, such as through a straddle. This concentrated demand raises the prices of these options, pushing up the implied volatilities at these strikes. Additionally, positive correlations between the payoffs at in-the-money and out-of-the-money strikes amplify this price impact across these strikes (e.g., as seen with strikes $x_{1}$ and $x_{3}$ in Figure 4). Since the implied volatility increases with option price, this cross-strike impact raises implied volatilities as strikes move further from being at-the-money, producing the characteristic smile shape. This mechanism explains how cross-strike price impacts can aggregate volatility information across options into a volatility smile.
Furthermore, as the true volatility increases, the smile becomes more convex (see Figure 8b), offering a testable empirical prediction. The underlying mechanism is that a higher volatility increases the covariance between payoffs at the in-the-money and out-of-the-money strikes, amplifying the cross-strike price impact and making the smile more pronounced.
9 Conclusion
By unifying the elements of Arrow and Debreu (1954) and Kyle (1985), we have developed a general model of informed trading across contingent claim markets. Our findings advance the theoretical understanding of how contingent claim markets jointly incorporate information. Our framework provides explanations for observed empirical practices, bridges the gap with empirical literature, and proposes new avenues for future research.
Appendix A Appendix
A.1 Model Assumptions
The formal assumptions of the general model are as follows.
Assumption 1.
(i) Possible signals lie in the finite probability space $(S,\pi_{0})$ where $S=\{s_{1},\cdots,s_{K}\}$, and the probability measure $\pi_{0}$ is the market maker’s prior.
(ii) Possible realizations order flow across states lie in the measurable space $(\Omega,\mathcal{F})$, where $\Omega=\mathbb{R}^{\infty}$ (the set of countable sequences) and $\mathcal{F}$ is the Borel $\sigma$-field generated by the coordinate functions
$\omega=(\omega_{i})\mapsto\omega_{j},\;\Omega\rightarrow\mathbb{R},\;\;\mbox{$j=1,2,\cdots$}.$
(iii) The probability measure $\mathbb{P}_{\!\scriptscriptstyle 0}$ on $(\Omega,\mathcal{F},\mathbb{P}_{\!\scriptscriptstyle 0})$ specifies the canonical process $\omega\mapsto\omega_{i}$, $i=1,2,\cdots$, as the stochastic sequence
$\varepsilon_{i}\!\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})\;\;\mbox{$i=1,2,\cdots$}.$
In other words, $\mathbb{P}_{\!\scriptscriptstyle 0}$ specifies the distribution of noise-only order flow $\omega$ when insider demand is zero across markets.
(iv) The insider’s AD portfolio $W\colon X\rightarrow\mathbb{R}$ satisfies
$\exp\left(\frac{1}{2}\sum\limits_{i=1}^{\infty}\left(\frac{W(x_{i})}{\sigma_{i}}\right)^{2}\right)<\infty.$
A.2 Proof of Theorem 4.1
Recall the following standard result (see (Shiryaev, 1999, Chapter V, Section 3)): Under Assumption 1(iv), for any insider portfolio $W\colon X\rightarrow\mathbb{R}$, consider the probability measure $\mathbb{P}_{\!{\scriptscriptstyle W}}$ on $(\Omega,\mathcal{F})$ defined by the Radon-Nikodym density
$\frac{\mathrm{d}\mathbb{P}_{\scriptscriptstyle W}}{\mathrm{d}\mathbb{P}_{\!\scriptscriptstyle 0}}=\exp\left(\sum\limits_{i=1}^{\infty}\frac{W(x_{i})\omega_{i}}{\sigma_{i}^{2}}-\frac{1}{2}\sum\limits_{i=1}^{\infty}\frac{W(x_{i})^{2}}{\sigma_{i}^{2}}\right).$ (A.1)
On $(\Omega,\mathcal{F},\mathbb{P}_{\!{\scriptscriptstyle W}})$, the canonical process $\omega\mapsto\omega_{i}$ specifies the stochastic sequence
$W(x_{i})+\varepsilon_{i},\;i=1,2,\cdots.$
So $\mathbb{P}_{\!\scriptscriptstyle W}$ specifies the distribution of order flow $\omega$ when insider demand is $W$.
Therefore, the probability measure on $\Omega\times S$ given by
$\left(\exp\left(\sum\limits_{i=1}^{\infty}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}-\frac{1}{2}\sum\limits_{i=1}^{\infty}\frac{\widetilde{W}(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}\right)\cdot\mathbb{P}_{\!\scriptscriptstyle 0}\right)\otimes\pi_{0}$
correctly specifies the intended joint probability law of $(\omega,s_{k})$ according to the market maker’s belief $\widetilde{W}(\,\cdot\,|\,\cdot\,)$. Since
$\exp\left(\sum\limits_{i=1}^{\infty}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}-\frac{1}{2}\sum\limits_{i=1}^{\infty}\frac{\widetilde{W}(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}\right)$
is clearly jointly measurable in $(\omega,s_{k})$, it follows from the Fubini-Tonelli Theorem that $\{\pi_{1}(\,\cdot\,|\omega\kern 0.44998pt;\widetilde{W})\}_{\omega\in\Omega}$ is the $\omega$-disintegration of the family $\{\mathbb{P}_{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}\}_{s_{k}\in S}$.
In other words, the posterior $\pi_{1}(\,\cdot\,|\omega\kern 0.44998pt;\widetilde{W})$ of (11) specifies the market maker’s posterior probability measure on $S$ conditional on $\omega$, with the normalizing random variable
$C(\omega)=\sum\limits_{k=1}^{K}\exp\left(\sum\limits_{i=1}^{\infty}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}-\frac{1}{2}\sum\limits_{i=1}^{\infty}\frac{\widetilde{W}(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}\right)\cdot\pi_{0}(s_{k}).$
This proves the theorem.
A.3 Proof of Theorem 5.1
The marginal utility functional $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)\colon\mathbb{R}^{\infty}\rightarrow\mathbb{R}$ is the Gâteaux-derivative of the expected utility functional $J(\,\cdot\,|s_{k};\widetilde{W})\colon\mathbb{R}^{\infty}\rightarrow{\mathbb{R}}$, as defined in Equation (14), at $W$. (Both $J(\,\cdot\,|s_{k};\widetilde{W})$ and its Gâteaux-derivative $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)$ are conditional on $s_{k}$ and parameterized by $\widetilde{W}$.) $J(W|s_{k};\widetilde{W})=J_{p}(W)-J_{c}(W)$ is the difference between
[…]
and
[…]
We need to show its Gâteaux-derivative decomposes into
$\underbrace{\partial\!\kern 0.80002ptJ(\,\cdot\,;W)}_{\text{marginal utility}}=\underbrace{\partial\!\kern 0.80002ptJ_{p}(\,\cdot\,;W)}_{\text{marginal payoff}}-\underbrace{\big{(}\,\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(\,\cdot\,;W)+\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(\,\cdot\,;W)\,\big{)}}_{\text{marginal cost}}$
Because the Gâteaux derivative of the payoff functional $J_{p}(\,\cdot\,)$ is trivially identified with $\eta(\,\cdot\,|s_{k})$, the characterization of the marginal payoff functional $\partial\!\kern 0.80002ptJ_{p}(\,\cdot\,;W)$ in Equation (15) is immediate.
Now, consider the cost functional $J_{c}(\,\cdot\,)$. To make the dependence of expected AD price $\overline{P}(x_{i},W;\widetilde{W})$ on $W$ more explicit, we write out
$\nopagebreak\overline{P}(x_{i},W;\widetilde{W})$ $=\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[\sum_{k}\eta(x_{i}|s_{k})\pi_{1}(s_{k},\omega;\widetilde{W})]$
$=\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle W}}[\sum_{k}\eta(x_{i}|s_{k})\,\frac{{\rm e}^{\mathcal{I}(\omega,s_{k};\widetilde{W})}}{C(\omega)}\,\pi_{0}(s_{k})]$
$=\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle 0}}[\sum_{k}\eta(x_{i}|s_{k})\,\frac{{\rm e}^{\mathcal{I}(\omega,s_{k};\widetilde{W})+\sum\limits_{j}\frac{\widetilde{W}(x_{j}|s_{k})W(x_{j})}{\sigma_{j}^{2}}}}{C^{\prime}(\omega)}\,\pi_{0}(s_{k})],$ (A.2)
where
$\mathcal{I}(\omega,s_{k};\widetilde{W})=\sum_{j}\frac{\widetilde{W}(x_{j}|s_{k})\omega_{j}}{\sigma_{j}^{2}}-\frac{1}{2}\sum_{j}\frac{\widetilde{W}(x_{j}|s_{k})^{2}}{\sigma_{j}^{2}}$ (A.3)
and
$C^{\prime}(\omega)=\sum_{l}{\rm e}^{\mathcal{I}(\omega,s_{l};\widetilde{W})+\sum_{j}\frac{\widetilde{W}(x_{j}|s_{l})W(x_{j})}{\sigma_{j}^{2}}}\pi_{0}(s_{l}).$ (A.4)
The equality (A.2) holds because the law of the canonical process $\omega\mapsto\omega_{j}$ under $\mathbb{P}_{\scriptscriptstyle W}$ is the same as the law of $\omega\mapsto W(x_{j})+\omega_{j}$ under $\mathbb{P}_{\!\scriptscriptstyle 0}$.
Let $v\colon X\rightarrow{\mathbb{R}}$ be a marginal portfolio and define $f(\varepsilon)=J_{c}(W+\varepsilon v)$. The marginal cost functional (i.e., Gâteaux derivative) $\partial\!\kern 0.80002ptJ_{c}(W)$ of $J_{c}(W)$ evaluated at $v$ can be computed by invoking the Dominated Convergence Theorem and differentiating under the summation signs:
$\partial\!\kern 0.80002ptJ_{c}(W)$ $=f^{\prime}(0)$
$=\sum_{i}v(x_{i})\overline{P}(x_{i},W;\widetilde{W})+\sum_{i}W(x_{i})g(x_{i})$ (A.5)
for some $g\colon X\rightarrow\mathbb{R}$. The first sum in (A.3) verifies the AD term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!AD}(v;W)$ of Equation (16). It remains to show the price impact term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ of Equation (17) is the second sum in (A.3).
The function $g$ is of the form $g(x_{i})=\mathbb{E}^{\mathbb{P}_{\!\scriptscriptstyle 0}}[\psi(\omega;x_{i},\widetilde{W})]$, $i=1,2,\cdots$. The random variable $\psi(\,\cdot\,;x_{i},\widetilde{W})$ defined on $\Omega$ is given by
$\psi(\omega;x_{i},\widetilde{W})=\sum_{k}\left(\eta(x_{i}|s_{k})\frac{C^{\prime}(\omega)l(s,\omega)\sum\limits_{j}\frac{\widetilde{W}(x_{j}|s_{k})}{\sigma_{j}^{2}}v(x_{j})-l(s,\omega)\sum\limits_{l}\left(l(s_{l},\omega)\sum\limits_{j}\frac{\widetilde{W}(x_{j}|s_{l})}{\sigma_{j}^{2}}v(x_{j})\pi_{0}(s_{l})\right)}{C^{\prime}(\omega)^{2}}\pi_{0}(s_{k})\right)$ (A.6)
where
$C^{\prime}(\omega)=\sum_{k}\left({\rm e}^{\sum\limits_{j}\frac{\widetilde{W}(x_{j}|s_{k})W(x_{j})}{\sigma_{j}^{2}}+\cdots}\pi_{0}(s_{k})\right)=\sum_{k}l(s_{k},\omega)\pi_{0}(s_{k})$
is the normalization constant of the market maker’s posterior under the probability measure $\mathbb{P}_{\!\scriptscriptstyle 0}$, and
$l(s_{k},\omega)={\rm e}^{\sum\limits_{j}\frac{\widetilde{W}(x_{j}|s_{k})W(x_{j})}{\sigma_{j}^{2}}+\cdots}$ (A.7)
is the likelihood of $\omega$ under $\mathbb{P}_{\!\scriptscriptstyle 0}$. For clarity of notation, in (A.7) we have put $\cdots$” for terms not relevant for this calculation.
Now, to intrepret $g(x_{i})=\mathbb{E}^{\mathbb{P}_{\!\scriptscriptstyle 0}}[\psi(\omega;x_{i},\widetilde{W})]$, observe that
$\frac{l(s_{k},\omega)\pi_{0}(s_{k})}{C^{\prime}(\omega)},\;k=1,\cdots,K$
is the market maker’s posterior condition on $\omega$. It is then clear from Equation (A.6) that $\psi(\omega;x_{i},\widetilde{W})$ is the difference between the posterior expectation of the product of $\eta(x_{i}|\,\cdot\,)$ and $\Pi_{insider}(v,\,\cdot\,;\widetilde{W})=\sum\limits_{j}\frac{\widetilde{W}(x_{j}|\,\cdot\,)v(x_{j})}{\sigma_{j}^{2}}$ (as in Definition 2) and the product of their posterior expectations. In other words, $\psi(\omega;x,\widetilde{W})$ is equal to the posterior covariance
$\psi(\omega;x,\widetilde{W})={\operatorname{\mathbb{Cov}}}\!\left(\eta(x,\,\cdot\,),\Pi_{insider}(v,\,\cdot\,;\widetilde{W})\big{|}\omega\right).$
This shows that the second term in $\partial J_{c}(W)$ is precisely the price impact term $\partial\!\kern 0.80002ptJ_{\scriptscriptstyle\!K}(v;W)$ of Equation (17). This proves the theorem.
A.4 Proof of Theorem 5.4
For $c\colon S\rightarrow{\mathbb{R}}$, define the affine subspace of portfolios $\mathcal{V}_{c}(\widetilde{W})\!\equiv\!\{W\colon\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=c\}.$ The zero price impact portfolios are those in the subspace $\mathcal{V}_{0}(\widetilde{W})$ corresponding to $c=0$.
We observe that, for a given $c$, the market maker’s (expected) AD prices $\overline{P}(\,\cdot\,,W;\widetilde{W})$ does not change with respect to $W\!\in\!\mathcal{V}_{c}(\widetilde{W})$. [^24] This can be seen explicitly from Equation (A.2). That is, conditional on $\Pi_{insider}(W,\,\cdot\,;\widetilde{W})$, a portfolio $W$ does not change the (expected) AD prices. In particular, this is true for the zero price impact subspace $\mathcal{V}_{0}(\widetilde{W})$. Also, $\mathcal{V}_{0}(\widetilde{W})$ is invariant under scaling. It follows that, for the insider’s problem (14) to be well-posed, a portfolio $W\in\mathcal{V}_{0}(\widetilde{W})$ must give him zero expected utility. Otherwise, he can obtain unbounded utility by scaling up his portfolio indefinitely without incurring price impact. For example, suppose $J(W|s_{k};\widetilde{W})>0$ for some $W\in\mathcal{V}_{0}(\widetilde{W})$. Then the insider’s expected utility $J(\alpha W;\widetilde{W},s)\rightarrow\infty$ as $\alpha\rightarrow\infty$ because the scaled portfolio $\alpha W\in\mathcal{V}_{0}(\widetilde{W})$ causes no price impact as $\alpha\rightarrow\infty$. Similarly, $J(W|s_{k};\widetilde{W})<0$ for some $W\in\mathcal{V}_{0}(\widetilde{W})$ would allow arbitrage.
More generally, for any two portfolios $W_{1},W_{2}\in\mathcal{V}_{c}(\widetilde{W})$, no-arbitrage requires that $J(W_{1}|s_{k};\widetilde{W})=J(W_{2}|s_{k};\widetilde{W})$. Otherwise, the long-short portfolio $W_{1}-W_{2}\in\mathcal{V}_{0}(\widetilde{W})$ would be an arbitrage opportunity. In other words, to preclude arbitrage, the insider’s expected utility functional $J(\,\cdot\,|s_{k};\widetilde{W})$ must be constant on the closed affine subspace $\mathcal{V}_{c}(\widetilde{W})$ for each $c\in C(S,{\mathbb{R}})$. This proves the theorem.
A.5 Proof of Theorem 6.2
Under the general specification, the spanning conditions of Proposition 6.1 imply the equilibrium restriction that the noise trading intensity is a constant $\sigma\!>\!0$ across markets. Under additional assumptions (e.g., when $\eta(\,\cdot\,|s_{k})$’s have disjoint supports as in Section 2), this spanning condition can hold with a varying noise intensity. In such cases, the proof here goes through verbatim with $\sigma_{i}$, $i=1,2,\cdots$, in place of $\sigma$.
For order flow $\omega\in\Omega$, let $B(\omega)=(\sum\limits_{i}\frac{\eta(x_{i}|s_{k})\omega_{i}}{\sigma^{2}})_{k=1,\cdots,K}\in\mathbb{R}^{K}$. Then, in terms of $d$ and $\widetilde{D}$ (the latter is defined in (23)), the general market maker posterior over signals obtained in Theorem 4.1(i) can be written explicitly as the probability mass function
$\pi_{1}(k|\omega,d;\widetilde{D})=\frac{e^{(\tilde{d}^{(k)})^{T}{\bf L}^{2}d+(\tilde{d}^{(k)})^{T}B(\omega)-\frac{1}{2}(\tilde{d}^{(k)})^{T}{\bf L^{2}}\tilde{d}^{(k)}}}{\sum\limits_{l=1}^{K}e^{(\tilde{d}^{(l)})^{T}{\bf L^{2}}d+(\tilde{d}^{(l)})^{T}B(\omega)-\frac{1}{2}(\tilde{d}^{(l)})^{T}{\bf L}^{2}\tilde{d}^{(l)}}},\;\;k=1,\cdots,K$ (A.8)
on $S$. For the expectation of $\big{(}\pi_{1}(k|\omega,d;\widetilde{D})\big{)}_{k=1,\cdots,K}\in\mathbb{R}^{K}$, we write it as
$\bar{\pi}_{1}(d;\widetilde{D})=\left(\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle 0}}[\pi_{1}(k|\omega,d;\widetilde{D})]\right)_{k=1,\cdots,K}\in\mathbb{R}^{K},$ (A.9)
where $\mathbb{E}^{\mathbb{P}_{\scriptscriptstyle 0}}[\,\cdot\,]$ is taken with respect to distribution over possible order flows $\omega$ if the insider chooses $d$ and the market maker holds belief $\widetilde{D}$. The insider’s portfolio choice problem (14) conditional on observing $s_{k}$ now takes the simple form
$\max_{d\in\mathbb{R}^{K}}e_{k}^{T}{\bf L}^{2}d-\bar{\pi}_{1}(d;\widetilde{D})^{T}{\bf L}^{2}d\equiv\max_{d\in\mathbb{R}^{K}}J(d|k;\widetilde{D}).$ (A.10)
This reduces the Bayesian trading game between the insider and the market maker to one where the market maker’s posterior belief is specified by (A.8), and the insider’s problem is (A.10).
The canonical transformation of (22)
$\hat{d}={\bf L}d$
replaces ${\bf L}^{2}$ and $d$ in the market maker’s posterior $\pi_{1}(k|\omega,d;\widetilde{D})$ of (A.8) by ${\bf I}$ and $\hat{d}$, respectively. Under this transformation, the $k$-th component of the random vector $\widehat{N}={\bf L}^{-1}B(\omega)$ can be re-written as $\widehat{N}_{k}=\sum\limits_{i}\xi^{(k)}_{i}\tilde{\varepsilon}_{i}$ where $\sum\limits_{i}\xi^{(k)}_{i}\xi^{(l)}_{i}=\delta_{kl}$ and $(\tilde{\varepsilon}_{i})$ is an i.i.d. standard normal sequence. Therefore, by (the discrete version of) Itô isometry,
$\widehat{N}_{k}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,1),\;\;k=1,\cdots,K.$ (A.11)
This gives the market maker’s posterior (24) in the canonical game.
Similarly, the canonical transformation replaces ${\bf L}^{2}$ and $d$ in the insider’s objective function $J(d|k;\widetilde{D})$ of (A.10) by ${\bf I}$ and $\hat{d}$, respectively. This verifies the insider’s problem (25) in the canonical game and proves the proposition.
A.6 Proof of Theorem 6.5
Lemma A.1.
Suppose $\alpha^{\prime}>0$ solves the equilibrium equation $\Phi(\alpha)=0$, and suppose the market maker holds belief $\alpha^{\prime}{\bf Q}$ in the canonical game. Then, for the insider’s problem (25) conditional on each $s_{k}$ in the canonical game, the first-order condition is sufficient for optimality. Therefore, since the insider’s strategy $\alpha^{\prime}{\bf Q}$ satisfies the collated first-order conditions $\Phi(\alpha^{\prime}){\bf Q}=0$, $\alpha^{\prime}{\bf Q}$ is an equilibrium of the canonical game as defined in (26).
Proof.
Given the market maker’s belief $\widehat{D}=\alpha^{\prime}{\bf Q}$ in the canonical game, the Hessian matrix $H$ of the insider’s objective function (25) conditional on a signal $s_{k}$ can be computed by differentiating directly his first-order condition (28), which gives
$-\alpha^{\prime}{\bf Q}\underbrace{\mathbb{E}\left[\begin{bmatrix}p_{1}&-p_{1}p_{2}&\cdots\\ -p_{1}p_{2}&p_{2}&\cdots\\ \vdots&\vdots&\ddots\end{bmatrix}\right]}_{\text{$H$}}=-\alpha^{\prime}{\bf Q}H.$
By the Cauchy-Schwarz inequality,
[…]
Therefore $H$ is positive semidefinite. Since ${\bf Q}H=H{\bf Q}$, ${\bf Q}H$ is also positive semidefinite. Therefore the Hessian $-\alpha^{\prime}{\bf Q}H$ is negative semidefinite. It follows that the insider’s objective function is concave conditional on any given signal. This proves the lemma. ∎
Lemma A.2.
Let
${\bf p}=[p_{1}\,\cdots\,p_{K}]^{T}$
denote the random vector (24) of the market maker’s (normalized) posterior in the canonical game conditional on the insider observing signal $s_{k}$ (with the dependence on $\hat{\omega}=\hat{d}+\widehat{N}$, and $\widehat{D}$ understood). ${\bf p}$ is a random probability measure on the signal space $S=\{s_{1},\cdots,s_{K}\}$. Let $\alpha>0$ and $Z\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,\alpha^{2}{\bf Q})$ be a random vector that is multivariate normal with mean 00 and covariance matrix $\alpha^{2}{\bf Q}$.
[…]
${\bf p}\stackrel{{\scriptstyle d}}{{\sim}}\left(\frac{e^{Z_{1}}}{\sum\limits_{l}\cdots},\cdots,\frac{e^{\alpha^{2}+Z_{k}}}{\sum\limits_{l}\cdots},\cdots\frac{e^{Z_{K}}}{\sum\limits_{l}\cdots}\right)^{T}.$
Proof.
Substituting the equilibrium postulate $\alpha{\bf Q}$ into the market maker’s posterior (A.8) conditional on order flow $\omega$, we have
$\pi_{1}(s_{k}|\omega,\beta^{(k)};\alpha{\bf Q})=\frac{e^{\alpha^{2}(1-\frac{1}{K})\,+\,\alpha e_{k}^{T}{\bf Q}{\bf L}^{-1}B(\omega)}}{\sum\limits_{l}\cdots},$ (A.12)
and, for $m\neq k$,
$\pi_{1}(s_{m}|\omega,\beta^{(k)};\alpha{\bf Q})=\frac{e^{\alpha^{2}(-\frac{1}{K})\,+\,\alpha e_{m}^{T}{\bf Q}{\bf L}^{-1}B(\omega)}}{\sum\limits_{l}\cdots},$ (A.13)
where $\beta^{(k)}$ is the $k$-th column of $\alpha{\bf Q}$, i.e., the insider strategy postulated by $\alpha{\bf Q}$ conditional on observing signal $s_{k}$, and the denominator $\sum\limits_{l}\cdots$ is the normalization factor.
The common factor $e^{\alpha^{2}(-\frac{1}{K})}$ in (A.12) and (A.13) cancels after normalization. It remains to consider the random vector ${\bf L}^{-1}B(\omega)$. By the same Itô isometry argument as that for (A.11), ${\bf L}^{-1}B(\omega)\in\mathbb{R}^{K}$ has the standard multivariate normal distribution. Therefore $\alpha{\bf Q}{\bf L}^{-1}B(\omega)\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,\alpha^{2}{\bf Q})$. This proves the lemma. ∎
Proof of Theorem
The Equilibrium Equation.
Substituting the equilibrium postulate $\widehat{D}=\alpha{\bf Q}$ of (29) into the insider’s first-order condition (28) for the canonical game conditional on (say) $s_{1}$ gives
$\mathbb{E}\left[e_{1}-\begin{bmatrix}p_{1}\\ \vdots\\ p_{K}\end{bmatrix}-\alpha^{2}{\bf Q}\begin{bmatrix}p_{1}&&\\ &\ddots&\\ &&p_{K}\end{bmatrix}{\bf Q}e_{1}+\alpha^{2}{\bf Q}\begin{bmatrix}p_{1}\\ \vdots\\ p_{K}\end{bmatrix}\begin{bmatrix}p_{1}&\cdots&p_{K}\end{bmatrix}{\bf Q}e_{1}\right]=0,$ (A.14)
where $\mathbb{E}[\,\cdot\,]$ is taken with respect to the probability law of the random probability measure ${\bf p}=\begin{bmatrix}p_{1}\\ \vdots\\ p_{K}\end{bmatrix}$ under the postulated equilibrium.
The first-order condition (A.14) conditional on $s_{1}$ simplifies to
$\mathbb{E}\left[e_{1}-\begin{bmatrix}p_{1}\\ p_{2}\\ \vdots\\ p_{K}\end{bmatrix}-\alpha^{2}\big{(}\begin{bmatrix}p_{1}\\ 0\\ \vdots\\ 0\end{bmatrix}-p_{1}\begin{bmatrix}p_{1}\\ p_{2}\\ \vdots\\ p_{K}\end{bmatrix}\big{)}\right]=\mathbb{E}\left[(1-\alpha^{2}p_{1})\begin{bmatrix}1-p_{1}\\ -p_{2}\\ \vdots\\ -p_{K}\end{bmatrix}\right]=0.$ (A.15)
The probability law of ${\bf p}$ is as characterized in Lemma A.2 above, with $k=1$. Under this probability law, the moments in Equation (A.15) are the moments of a logistic-normal distribution.
We now show that, by substituting for the appropriate relationships between corresponding moments, (A.15) can be written as
$\Phi(\alpha){\bf Q}e_{1}=0,$ (A.16)
for some $\Phi\colon[0,\infty)\rightarrow\mathbb{R}$. To show this, it suffices to show
$\mathbb{E}[(1-\alpha^{2}p_{1})(1-p_{1})]=(K-1)\mathbb{E}[(1-\alpha^{2}p_{1})p_{l}],\;\;l=2,\cdots,K.$ (A.17)
(A.17) reduces to
$\mathbb{E}[(1-\alpha^{2}p_{1})(1-p_{1})]=(K-1)\mathbb{E}[(1-\alpha^{2}p_{1})p_{2}]$ (A.18)
because, under the probability law of ${\bf p}$ conditional on $s_{1}$ characterized in Lemma A.2, $\mathbb{E}[p_{1}p_{l}]=\mathbb{E}\left[p_{1}p_{m}\right]$ for all $l,m\neq 1$. In turn, (A.18) holds because
$\mathbb{E}[1-p_{1}]=(K-1)\mathbb{E}[p_{2}],\,\mbox{and}\,\,\,\mathbb{E}[p_{1}(1-p_{1})]=\mathbb{E}[p_{1}(p_{2}+\cdots+p_{K})]=(K-1)\mathbb{E}[p_{1}p_{2}]$
under the same logistic-normal probability law.
Therefore, we can take $\Phi(\alpha)$ in (A.16) to be (up to a scalar multiple) the left-hand side of (A.17), i.e.,
$\Phi(\alpha)=\mathbb{E}[1-p_{1}-\alpha^{2}p_{1}+\alpha p_{1}^{2}]$
where
$p_{1}=\frac{e^{\alpha^{2}+Z_{1}}}{e^{\alpha^{2}+Z_{1}}+\sum\limits_{l\neq 1}e^{Z_{l}}},\;\;Z=(Z_{l})_{1\leq l\leq K}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,\alpha^{2}{\bf Q}).$
By symmetry, the first-order condition for $k\neq 1$ is identical to (A.16) after permuting the indices $1$ and $k$. Collating these $K$ symmetric first-order conditions,
[…]
side-by-side gives the matrix equation $\Phi(\alpha){\bf Q}=0$ of (30). By Lemma A.1, an equilibrium is given by a solution $\alpha^{*}>0$ to the equilibrium equation $\Phi(\alpha)=0$.
Existence of Equilibrium.
First, we have $\Phi(0)=\mathbb{E}[1-p_{1}]>0$. Second, $1-p_{1}-\alpha^{2}p_{1}+\alpha p_{1}^{2}\rightarrow 0$ and $01-p_{1}-\alpha^{2}p_{1}+\alpha p_{1}^{2}<$ 0 eventually as $\alpha\rightarrow\infty$, with probability one. By Fatou’s Lemma, we have that $\Phi(\alpha)\rightarrow 0$ from below as $\alpha\rightarrow\infty$. Therefore, by the Intermediate Value Theorem, there exists $\alpha^{*}>0$ such that $\Phi(\alpha^{*})=0$. This proves the theorem.
A.7 Proof of Proposition 7.7
This is a special case of Lemma A.2, with $\alpha=\alpha^{*}$.